### v1

# Adapting Pre-Trained Vision Models for Novel Instance Detection and Segmentation

Abstract
Novel Instance Detection and Segmentation (NIDS) aims at detecting and segmenting novel object instances given a few examples of each instance. We propose a unified framework (NIDS-Net) comprising object proposal generation, embedding creation for both instance templates and proposal regions, and embedding matching for instance label assignment. Leveraging recent advancements in large vision methods, we utilize the Grounding DINO and Segment Anything Model (SAM) to obtain object proposals with accurate bounding boxes and masks. Central to our approach is the generation of high-quality instance embeddings. We utilize foreground feature averages of patch embeddings from the DINOv2 ViT backbone, followed by refinement through a weight adapter mechanism that we introduce. We show experimentally that our weight adapter can adjust the embeddings locally within their feature space and effectively limit overfitting. This methodology enables a straightforward matching strategy, resulting in significant performance gains. Our framework surpasses current state-of-the-art methods, demonstrating notable improvements of 22.3, 46.2, 10.3, and 24.0 in average precision (AP) across four detection datasets. In instance segmentation tasks on seven core datasets of the BOP challenge, our method outperforms the top RGB methods by 3.6 AP and remains competitive with the best RGB-D method [^1] Code is available at: https://github.com/YoungSean/NIDS-Net.
1 Introduction
Novel Instance Detection and Segmentation (NIDS) is a crucial task in computer vision aimed at identifying and locating unseen instances in images or videos, given a few examples of each instance. Of particular interest in this work, are applications of NIDS in robotics, e.g., a robot might need to fetch a specific, novel object instance from a cluttered desk given only a small number of multi-view template images of the novel object. Ideally, detectors for this task should provide precise bounding boxes and masks of the target instances in a cluttered query image, but standard detection systems, trained on specific classes, often struggle to recognize new objects.
The current paradigm for NIDS typically encompasses the following steps: (1) generating proposals from a query image, (2) obtaining embeddings of the proposals and the target instance templates, and (3) matching the embeddings of proposals with those of the templates for identification. Recent work [45, 79, 64, 49] has utilized various open-world detectors to obtain object proposals. However, the use of models such as the Segment Anything Model (SAM) [43] or FastSAM [101] often results in the generation of region-based proposals rather than actual object proposals. This misidentification of regions as objects can to lead to object identification issues. For example, a single object may be divided into multiple proposals or some background regions may be mistaken for foreground objects. These false alarms not only slow down the whole detection progress but also spoil the final results. To represent proposals and templates, Shen et al. [79], Nguyen et al. [64], and Lin et al. [49] adopt the $cls$ token of DINOv2  [66, 20] while Li et al. [45] employ a 3D voxel representation. Ideally, for each specific unseen instance, the template embeddings from different views should be similar to each other but markedly different from the embeddings of other instances. Despite these efforts, the challenge persists that the embeddings of different instances remain excessively similar. The development of robust feature embeddings could provide a solution to this issue.
To address the limitations of current methods, we propose a framework for NIDS. Initially, we utilize Grounding DINO [52] with the text prompt “objects” on a cluttered query image to obtain bounding boxes for foreground objects. This method capitalizes on the inherent objectness of Grounding DINO using specialized prompts rather than standard labels. Subsequently, we employ SAM to create masks within these bounding boxes, generating precise object proposals comprising both bounding boxes and masks. For instance embeddings, we first extract average foreground features from the patch embeddings of the DINOv2 ViT backbone. We then apply adapters to enhance instance embeddings by clustering similar instances and distancing different ones. The adapters are trained with instance template embeddings via InfoNCE loss [65, 15]. To refine these embeddings, the CLIP-Adapter [28] introduces a residual vector added to the original embedding, which risks overfitting with a few training examples. This addition can spoil and destabilize the embeddings of background objects, disrupting the entire feature space. To mitigate this issue, we introduce a weight adapter that modifies the original embeddings by applying learned weights. Finally, we employ a straightforward matching approach [60], such as stable matching or using $argmax$, to assign instance labels to these proposals.
Our approach has been validated on four detection datasets and seven segmentation datasets. The framework significantly surpasses existing state-of-the-art methods: it demonstrates substantial increases in average precision (AP), with gains of 22.3, 46.2, 10.3, and 24.0 across these detection datasets. Moreover, in instance segmentation tasks on the seven core datasets of the BOP challenge [82], our method outperforms the best RGB and RGB-D methods.
Our contributions are summarized as follows.
  • We propose NIDS-Net, a unified framework for novel instance detection and segmentation that surpasses existing methods across four detection datasets and seven segmentation datasets of BOP challenge.
  • We leverage the objectness of open-world detector Grounding DINO and region recognition of Segmentation Anything Model (SAM) to obtain object proposals.
  • We introduce the Weight Adapter, a tool designed to refine embeddings within their feature space while discouraging overfitting. This approach yields increasingly significant improvements as the adapting of the feature space becomes more robust and stable.
2 Related Work
Pretrained models. Large-scale pretraining has demonstrated utility across various downstream tasks. Works such as [32, 90, 55, 24, 54, 10, 66] emphasize large-scale image representation learning, while [72, 46] focus on vision-language pretraining. Additionally, [3, 4, 95] address depth prediction, and [70, 71, 6, 84] tackle large-scale language modeling. Object detection is addressed by [30, 77, 75, 40, 41, 31, 52], while image segmentation is the focus of [33, 86, 42]. Works such as [93, 48, 14, 74] explore text-to-3D generation, and [80, 1] address text-to-video synthesis. Finally, 6DoF pose estimation and tracking are tackled by [87, 21, 47, 85]. These foundational models offer diverse capabilities, with the challenge lying in effectively leveraging their wealth of knowledge for specific use cases, such as novel instance detection. DINOv2 [66, 20] offers robust features to represent unseen instances.
Open-world object detection (OWOD). OWOD tackles real-world scenarios where object detection algorithms are tasked not only with identifying known and labeled objects but also with effectively handling novel and unknown objects absent during inference. This sets OWOD apart from traditional object detection, which confines itself to detecting only predetermined object classes. A key challenge in OWOD lies in detecting and classifying unknown objects that were not part of the training data. Recent approaches such as OWL-ViT [63], DITO [40], RO-ViT [41], YOLO-World [18], Sambor [31], and Grounding DINO [52] offer diverse solutions to address OWOD challenges. OWL-ViT leverages a standard Vision Transformer [24] architecture with minimal modifications, incorporating contrastive image-text pre-training and end-to-end detection fine-tuning. DITO introduces an open-vocabulary detection approach based on detection-oriented image-text pretraining, replacing the classification architecture with a detector architecture during pretraining to better serve region-level recognition needs. RO-ViT proposes cropping and resizing regions of positional embeddings (PE) at pretraining, aligning better with region-level detection fine-tuning. YOLO-World enhances YOLO [75] with open vocabulary detection capabilities through vision-language modeling and large-scale pretraining. Grounding DINO [52] utilizes features from DINO [10] to ground text queries, while Sambor extends SAM [42] to detect arbitrary objects based on human inputs such as category names or reference expressions. We employ Grounding DINO due to its strong performance.
Few-shot object detection (FSOD). FSOD involves detecting occurrences of a novel class, unseen during training, given only a few template examples. De-ViT [100] introduces a region-propagation mechanism for localization, transforming propagated region masks into bounding boxes using a learnable spatial integral layer. FS-DETR [7], the first few-shot detection transformer, extends DETR [9] to process arbitrary numbers of novel objects concurrently while supporting various examples from each class. Other approaches such as [27] prioritize salient representations through angle distance and magnitude dispersion across few-shot samples,  [92] leverages variational autoencoders for data augmentation, and  [59] learns geometry-aware features with an offline simplex equiangular tight frame classifier. Additionally,  [56] generates specific and representative prototypes for each query image, considering both support and query image features.
Instance detection. Instance detection involves identifying an unseen instance in a test image using corresponding templates. Some methods, e.g.,  [62, 2, 53], rely on pure 2D representations and matching techniques. Specifically, DTOID [62] employs global object attention and a local pose-specific branch for template-guided heatmap prediction. However, these methods may struggle with variations in 2D appearance caused by occlusion or pose changes. In contrast, VoxDet [45] utilizes explicit 3D knowledge from multi-view templates, providing geometry-invariant representations. We generate 2D robust instance embeddings from templates with DINOv2 [66, 20].
Unseen object instance segmentation (UOIS). UOIS is a task designed to segment unseen objects in cluttered scenes. UOIS generates object masks without providing labels. UOIS-Net [89] utilizes a two-stage framework, initially generating rough masks using depth information, and then refining them into precise instance masks with RGB data. Both UCN [88] and MSMFormer [57, 58] segment objects by clustering pixel embeddings. Although these methods can generate object proposals, they primarily focus on tabletop settings. In contrast, our approach is designed to handle RGB query images in various environments.
Adapter. The application of adapters atop large pre-trained models has emerged as a prominent strategy, yielding significant improvements across various tasks. Previous works have leveraged adapters to enhance few-shot image classification tasks [28, 97, 68, 99, 81]. Other works have demonstrated effectiveness in language model fine-tuning [36, 37, 98, 29]. Additional works exhibit the promise of adapters in vision tasks including detection and segmentation [17, 16, 69, 91, 23]. For novel instance detection and segmentation, we propose the Weight Adapter to enhance performance. The adapter assigns weights to the embeddings from frozen backbones.
3 Method
Novel instance detection and segmentation (NIDS), also referred to as open-world detection and segmentation, is a task to locate and label unseen instances within a query image. We will assume that each of the $N$ target instances is represented by $K$ template images $I_{T}\in\mathbb{R}^{K\times 3\times W\times H}$ and their corresponding segmentation masks. During inference, the output of each query image $I_{Q}\in\mathbb{R}^{3\times W\times H}$ provides these instances’ bounding boxes for detection and instance masks for segmentation.
3.1 Instance Embedding Generation Stage
In our approach, an instance embedding summarizes the pixels of an instance. The objective of this stage is to generate initial instance embeddings. Each of the $N$ instances has $K$ multi-view template images and their ground truth segmentation masks from which we derive template embeddings $E_{T}\in\mathbb{R}^{N\times K\times C}$, where $C$ is the dimension of the embeddings. Given an image and its corresponding mask, we initially extract patch embeddings using the ViT backbone of DINOv2, and subsequently obtain foreground features as specified by the mask. We then perform average pooling on these features. This process, termed Foreground Feature Averaging (FFA), is proposed by [44] to assess object similarity. We employ FFA to generate all initial instance embeddings.
3.2 Object Proposal Stage
The purpose of this stage is to acquire object proposals from a query image. Previous studies [79, 64, 49] have utilized SAM [43] or FastSAM [101] to generate regions as object proposals. Such proposals contain a high number of false alarms. For example, SAM might misclassify background regions or parts of objects as complete objects. To address this challenge, an off-the-shelf zero-shot detector, Grounding Dino [52], is employed with the text prompt “objects” to obtain initial bounding boxes of foreground objects. Then, SAM is applied to create masks based on these bounding boxes. The integration of these two models, termed Grounded-SAM (GS) [78], significantly reduces the number of erroneous object proposals and expedites the subsequent stages. Moreover, this approach eliminates the need for training of the detector. We then extract the proposal regions along with their corresponding masks from the query images. Using the FFA pipeline with proposals, we calculate proposal embeddings $E_{P}\in\mathbb{R}^{Q\times C}$, which correspond to $Q$ regions of interest. Fig. 1 illustrates the process of obtaining proposal embeddings.
Figure 1: In our framework NIDS-Net, only adapters are learnable, while other models are frozen. Instance IDs are the instance labels.
3.3 Embedding Refinement via an Adapter
Most instance embeddings are well separated. However, some embeddings from distinct instances may cluster together. For exxample, background object embeddings might be similar to targets. Some target instances may resemble each other. To address this issue, we employ learnable adapters to refine the embeddings. We train these adapters using InfoNCE loss [65, 15], aiming to bring the embeddings of the same instance closer together while separating the embeddings of different instances.
CLIP-Adapter (CA). CLIP-Adapter [28] comprises two trainable linear bottleneck layers appended to the language and image branches of CLIP. During few-shot fine-tuning, the original CLIP backbone remains frozen. Nonetheless, adding extra layers for fine-tuning may result in overfitting to the few examples available. To mitigate this issue, CLIP-Adapter integrates residual connections that dynamically merge the newly adapted knowledge with the foundational knowledge from the original CLIP backbone.
Given an instance embedding $f$ and a $MLP$ (Multi-Layer Perceptron), the CLIP-Adapter modifies the embedding according to the formula $f_{c}=\alpha\times MLP(f)+(1-\alpha)\times f$. In this equation, $f_{c}$ represents the adapted embedding, which is a linear combination of the transformed embedding $MLP(f)$ and the original embedding $f$. The residual ratio $\alpha$ is set as 0.6.
Our Weight Adapter (WA). The CLIP-Adapter refines instance embeddings by adding a new feature vector. However, this new vector might not align with the feature space of original instance embeddings, which can lead to overfitting. Since only several hundred template embeddings of target instances are trained and refined in some datasets, the embeddings of background objects can be spoiled as the process disrupts the distances between target and background instance embeddings. Assuming the feature space of original embeddings is robust and effective, it is crucial to fine-tune the instance embeddings within this space by constraining the scope of the adapted embeddings. This refinement can be accomplished by simply applying weights to the original instance embeddings.
Hence, we introduce our Weight Adapter, a compact MLP network depicted in Fig. 2. The operation of this adapter is described through the following equations:
$w$ $=\text{sigmoid}(MLP(\beta f)),$ (1)
$f_{w}$ $=w\odot(\beta f).$ (2)
Here, $\beta$ is a positive scalar. We set $\beta$ as 10. $w$ is the weights computed by passing the scaled embedding $\beta f$ through the MLP and then applying a $sigmoid$ function. The output $f_{w}$ is the element-wise product of $w$ and the scaled embedding $\beta f$, resulting in the weighted embedding.
We include a ReLU layer before the $sigmoid$ function, ensuring that the sigmoid’s output is constrained to the range $[0.5,1)$. Subsequently, each refined embedding is close to its original counterpart. Additionally, $\beta$ allows for broader adjustments within the feature space. Given that cosine similarity is utilized to compare embeddings in the subsequent phase, our adapter alters the interrelationships among the values within each embedding. Further details of adapters are presented in Fig. 2. Qualitative results can be found in the appendix.
Figure 2: Comparison of two adapters. (left) CLIP-Adapter. (right) Our weight adapter.
3.4 Matching Stage
This stage provides each proposal $i$ with an instance ID $o_{i}$ as its label and a confidence score $s_{i}$.
Initially, $Q$ proposal embeddings $E_{P}$ are matched with $N\times K$ template embeddings $E_{T}$ using cosine similarity. It yields a matrix of template scores with dimensions $Q\times N\times K$, as illustrated in Fig. 1. For each instance, we aggregate all $K$ template scores to derive a matrix of instance scores with dimensions $Q\times N$. We employ $Max$ as the aggregation function for optimal results.
Bonus instance score for segmentation. To improve segmentation performance, we incorporate an additional appearance matching score, $s_{\text{appe}}$, as proposed by SAM-6D [49], into the instance scores. $s_{\text{appe}}$ is used to identify objects that are semantically similar yet differ in appearance. For each proposal, we can identify its most similar template $T_{\text{best}}$ according to template scores. $s_{\text{appe}}$ is derived from the patch embeddings of a proposal image $I$ and $T_{\text{best}}$. It quantifies the similarity between the query image and the template in terms of their local features as follows:
$s_{\text{appe}}=\frac{1}{N_{\text{patch}}^{I}}\sum_{j=1}^{N_{\text{patch}}^{I}}\max_{i=1,\ldots,N_{\text{patch}}^{T_{\text{best}}}}\left(CosineSimilarity(f_{I,j}^{\text{patch}},f_{T_{\text{best}},i}^{\text{patch}})\right)$
Assuming that objects are unique within a query image [79], we employ the stable matching algorithm [60] on instance scores to assign a unique instance ID to each proposal. If the assumption of uniqueness is not met, we use the $Argmax$ function on the instance score matrix, which permits multiple proposals to share the same instance ID.
After matching, we acquire labeled proposals. Each is defined by $\{b_{i},M_{i},o_{i},s_{i}\}$. Here, $b_{i}$ is the bounding box of an instance. $M_{i}$ is the modal mask, which covers the visible object surface [82]. $o_{i}$ is the instance ID. $s_{i}$ is the confidence score. A confidence score threshold $\delta$ can be used to remove incorrect proposals.
4 Experiments
We employ the Grounding DINO model with a Swin-T backbone [52] and a text prompt “objects”, and the default ViT-H SAM [43] to generate object proposals. According to the proposals, we obtain the regions of interest (RoIs) and resize them to $448\times 448$ or $224\times 224$ resolutions. For object detection, instance embeddings are produced using the DINOv2’s ViT-L model with registers [20], employing stable matching due to the uniqueness of instances in these datasets. For image segmentation, following SAM-6D [49], we utilize the ViT-L model of DINOv2 [66]. In scenarios with identical instances, such as the cluttered scenes in BOP datasets, we apply the $argmax$ function for matching.
The experiments are conducted on an RTX A5000 GPU. For each dataset, we train both adapters on template embeddings using InfoNCE loss [65, 15] with temperature 0.05. With two linear layers, Both adapters initially decrease the embedding dimension from $C$ to $C/4$, subsequently restoring it to $C$. The Weight Adapter is trained with Adam optimizer at a learning rate of 1e-3 and a batch size of 1024, while the CLIP-Adapter is trained at a learning rate of 1e-4. Dropout is incorporated in the CLIP-Adapter to reduce overfitting. For additional details, please see the Appendix.
4.1 Detection Datasets
We utilize recently developed instance detection datasets and their associated baselines [79, 45]. The High-resolution and RoboTools datasets employ real template images, while the LM-O and YCB-V datasets utilize synthetic images.
High-resolution dataset Shen et al. [79] design a high-resolution dataset for instance detection. This dataset comprises 100 distinct object instances, each represented by 24 photos taken from multiple viewpoints, with each photo having a resolution of $3072\times 3072$ pixels. These instances are integrated into 14 different indoor scenes for testing, captured in even higher resolution ($6144\times 8192$ pixels). The testing images are further classified as easy or hard based on the level of scene clutter and how much the instances are occluded. Easy tags are assigned when objects are sparsely placed, while hard tags are used for more cluttered setups.
Baselines According to Shen et al. [79], two types of baselines are set for the high-resolution dataset: Cut-Paste-Learn and a non-learned method. Cut-Paste-Learn generates synthetic training images with 2D-box annotations, putting foreground instances in various sizes and aspect ratios onto different backgrounds. This enables training detectors by viewing each instance as a class. They train five detectors: FasterRCNN [76], RetinaNet [51], CenterNet [102], FCOS [83], and the transformer-based DINO [96]. For the non-learned method, SAM is initially used to generate proposals, followed by employing DINO [11] and DINOv2 [66] to generate features for both proposal and template images, ultimately performing proposal matching and selection.
Synthetic-Real test sets Li et al. [45] employ two benchmarks for evaluation. LineMod-Occlusion (LM-O) [5] includes 8 texture-less objects and 1,514 bounding boxes, primarily challenging due to significant occlusion. The YCB-Video (YCB-V) [8] features 21 objects and 4,125 bounding boxes, with pose variation as the main challenge. Since these datasets have real testing images without reference videos, Li et al. [45] generate synthetic videos using CAD models. We sample 16 synthetic template images per object from these videos.
RoboTools benchmark Li et al. [45] introduce RoboTools, a challenging real-world benchmark for robotics. It features 20 distinct instances, 9,109 annotations, and 24 complex scenarios. Compared to previous benchmarks [5, 8], RoboTools offers greater difficulty due to its cluttered backgrounds and significant pose variations. Additionally, the benchmark’s reference videos are made from real images and incorporate realistic lighting conditions, including shadows. 25 real template images of each instance are sampled from reference videos.
Baselines For cluttered scenes from the datasets RoboTools, LM-O and YCB-V, Li et al. [45] have developed several 2D baselines: OLNDINO, OLNCLIP, and OLNCorr. Initially, they generate open-world 2D proposals using their detection module [39]. For matching, different methods are employed to select the proposal with the highest score. In OLNDINO and OLNCLIP, they use robust features from pre-trained backbones [11, 73] and cosine similarity for matching. For OLNCorr, a matching head is designed based on correlation [53]. They also employ the class-level one-shot detectors OS2D [67] and BHRL [94]. To address the limitations of traditional 2D methods with pose variations and occlusions, Li et al. [45] introduced VoxDet which is based on a 3D voxel representation. To train these methods, Li et al. [45] developed the open-world instance detection (OWID-10k) dataset as a synthetic instance detection training set. It features 9,901 objects from ShapeNet [12] and ABO [19], each rendered into a 40-frame, 360-degree video using Blenderproc [22].
Evaluation metrics For detection, we evaluate our method with Average Precision (AP). AP is computed by averaging precision scores at various Intersection over Union (IoU) thresholds, specifically from 0.5 to 0.95, in increments of 0.05. This metric is the primary evaluation standard in the COCO Object Detection dataset [50]. Additionally, AP50 and AP75 are variations of this metric, where precision is averaged across all instances at IoU thresholds of 0.5 and 0.75, respectively.
4.2 Segmentation Datasets
BOP challenge We test our method using seven datasets from the BOP challenge [82]: LineMod Occlusion (LM-O) [5], T-LESS [34], TUD-L [35], IC-BIN [25], ITODD [26], HomebrewedDB (HB) [38], and YCB-Video [8]. The objects from these datasets vary widely. We use 42 template rendering images from CNOS [64], which are generated via BlenderProc [22].
Baselines We compare our method with ZeroPose [13], CNOS [64], and SAM-6D [49]. These methods utilize proposals from SAM [43] or FastSAM [101] and use the $cls$ token from DINOv2 as the instance embedding for matching. In addition to adjusting the SAM hyperparameters to generate more proposals, SAM-6D enhances its performance by incorporating an appearance score, $s_{\text{appe}}$. Moreover, SAM-6D employs a Geometric Matching Score, $s_{\text{geo}}$, which considers the shapes and sizes of instances during matching by utilizing depth information.
Evaluation metrics For instance segmentation task, we evaluate our method using the Average Precision (AP) metrics. The AP is computed by averaging the precision scores at various IoU thresholds, from 0.50 to 0.95, increasing by 0.05 at each step.
Figure 3: Visual examples of our results using the Weight Adapter on the high-resolution and RoboTools datasets. Our approach detects specific object instances in cluttered scenes according to their real template images.
Table 1: Detection results on the high-resolution dataset. avg indicates that we are evaluating all images, including those from the easy and hard scenes. We train these two adapters with 2400 template embeddings.
Method AP AP50 AP75 Time (sec)
avg hard easy small medium large
FasterRCNN [76] 19.5 10.3 23.8 5.0 22.2 38.0 29.2 23.3 0.00399
RetinaNet [51] 22.2 14.9 26.5 5.5 25.8 42.7 31.2 25.0 0.00412
CenterNet [102] 21.1 11.8 25.7 5.9 24.1 40.4 32.7 23.6 0.00376
FCOS [83] 22.4 13.2 28.7 6.2 26.5 38.1 32.8 25.5 0.00271
DINO [96] 28.0 17.9 32.6 11.5 30.7 45.1 32.2 32.2 1.90625
SAM + DINOf [11] 37.0 22.4 43.9 11.9 40.9 62.7 44.1 40.4 15.10
SAM + DINOv2f [66] 41.6 28.0 47.6 14.6 45.8 69.1 49.1 46.0 14.70
NIDS-Net (Ours) 59.3 41.7 67.4 12.7 58.1 78.8 71.1 65.1 6.92
NIDS-Net + CA (Ours) 61.3 42.5 69.6 17.0 59.3 81.6 73.2 67.0 6.99
NIDS-Net + WA (Ours) 63.9 43.4 72.7 18.1 62.5 84.0 76.6 70.6 6.78
Table 2: Detection performance on the fully real dataset, RoboTools. Proposal indicates the object proposal model. GS stands for Grounded-SAM. OLN* is trained with the matching head while OLN employs fixed modules.
Method Proposal AP AP50 AP75
OS2D [67] N/A 2.9 6.5 2.0
DTOID [61] N/A 3.6 9.0 2.0
OLNCorr. [39, 53] OLN* 14.4 18.1 15.7
VoxDet [45] OLN* 18.7 23.6 20.5
NIDS-Net (Ours) GS 63.3 77.2 68.7
NIDS-Net + CA (Ours) GS 64.8 79.2 70.7
NIDS-Net + WA (Ours) GS 64.9 79.4 70.8
Table 3: Detection performance on the LM-O and YCB-V datasets. OLN* is trained alongside the matching head whereas OLN uses fixed modules. †means the model is trained on OWID and real images. Our running time of YCB-V is based on the scene 48. ()($\cdot$)( ⋅ ) represents the number of instances.
LM-O (8) YCB-V (21) Average
Method Proposal Train AP AP50 AP75 AP AP50 AP75 AP AP50 AP75 Time (s)
OS2D [67] N/A OWID 0.2 0.7 <0.1 5.2 18.3 1.9 2.7 9.5 1.0 0.189
DTOID [61] N/A OWID 9.8 28.9 3.7 16.3 48.8 42 13.1 38.9 4.0 0.357
OLNCLIP [39, 73] OLN OWID† 16.2 32.1 15.3 10.7 25.4 7.3 13.5 28.8 11.3 0.357
Gen6D [53] N/A OWID† 12.0 29.8 6.6 8.1 33.0 5.5 18.4 33.5 5.9 0.769
BHRL [94] N/A COCO 14.1 21.0 15.7 31.8 47.0 34.8 23.0 34.0 25.3 N/A
OLNCorr. [39, 53] OLN* OWID 22.3 34.4 24.7 24.8 41.1 26.1 23.6 37.8 25.4 0.182
OLNDino [39, 11] OLN OWID† 23.6 41.6 24.8 25.6 53.0 21.1 24.6 47.3 23.0 0.357
VoxDet [45] OLN* OWID 29.2 43.1 33.3 31.5 51.3 33.4 30.4 47.2 33.4 0.154
NIDS-Net (Ours) GS N/A 38.7 66.0 41.0 53.0 72.9 61.7 45.9 69.5 51.4 3.73
NIDS-Net + CA (Ours) GS N/A 39.2 67.0 41.4 53.9 74.1 62.7 46.6 70.6 52.1 3.71
NIDS-Net + WA (Ours) GS N/A 39.5 67.4 41.8 55.5 75.5 65.0 47.5 71.5 53.4 3.73
Table 4: Novel instance segmentation results on the seven core datasets of the BOP benchmark. We utilize Average Precision (AP) to compare these methods. ()($\cdot$)( ⋅ ) includes the number of instances.
Method BOP Datasets
LM-O (8) T-LESS (30) TUD-L (3) IC-BIN (2) ITODD (28) HB (33) YCB-V (21) Mean
ZeroPose (SAM) [13] 34.4 32.7 41.4 25.1 22.4 47.8 51.9 36.5
CNOS (SAM) [64] 39.6 39.7 39.1 28.4 28.2 48.0 59.5 40.4
CNOS (FastSAM) [64] 39.7 37.4 48.0 27.0 25.4 51.1 59.9 41.2
SAM-6D (FastSAM) [49] 39.5 37.6 48.7 25.7 25.3 51.2 60.2 41.2
SAM-6D (FastSAM) + $s_{\text{appe}}$ [49] 40.6 39.3 50.1 27.7 29.0 52.2 60.6 42.8
SAM-6D (SAM) [49] 43.4 39.1 48.2 33.3 28.8 55.1 60.3 44.0
SAM-6D (SAM) +$s_{\text{appe}}$ [49] 44.4 40.8 49.8 34.5 30.0 55.7 59.5 45.0
SAM-6D (SAM, RGBD[49] 46.0 45.1 56.9 35.7 33.2 59.3 60.5 48.1
NIDS-Net (Ours) 42.9 43.0 52.0 30.5 28.8 56.6 59.7 44.8
NIDS-Net + CA (Ours) 42.8 47.5 50.0 29.7 27.6 62.7 63.5 46.3
NIDS-Net + WA(Ours) 43.2 49.2 53.5 30.5 29.4 62.4 64.3 47.5
NIDS-Net + WA (Ours) + $s_{\text{appe}}$ 43.9 49.6 55.6 32.8 31.5 62.0 65.0 48.6
4.3 Benchmarking Results
Detection The high-resolution dataset results are presented in Table 1. Our models dramatically outperform existing techniques, achieving the highest AP scores across all categories. Our basic method surpasses the top baseline by 17.7 AP. Additionally, as this dataset contains 100 instances with some similarities among them, our Weight Adapter boosts overall performance by 4.6 AP and improves the detection of small objects by 5.4 AP. For the RoboTools dataset, detailed in Table 2, our method outperforms the state-of-the-art VoxDet by 46.2 AP. Since RoboTools contains only 20 instances, our adapter has limited scope to enhance performance. Our approach achieves over 60 AP on these two fully real datasets.
For the Synthetic-Real datasets, LM-O and YCB-V, we present a comparison of our approach with other methods in Table 3. Our model outperforms VoxDet by 10.3 AP on LM-O and by 24.0 AP on YCB-V. The improved performance on YCB-V, which contains more instances, indicates that our adapter functions more effectively with increased instance variety.
Segmentation CNOS [64] and SAM-6D [49] initially utilize SAM or FastSAM to obtain proposals and derive instance embeddings through the $cls$ token of Dinov2. In contrast, we generate proposals using Grounded-SAM and acquire embeddings through the FFA pipeline. The segmentation results for the seven principal datasets of the BOP challenge are detailed in Table 4. Overall, our method surpasses the state-of-the-art SAM-6D and achieves superior results compared to RGBD results.
Qualitative result We present detection results for both simple and complex scenes in Fig. 3. Additional detection and segmentation examples are available in the Appendix.
Runtime The tables of detection results include the runtime of our method. Despite Grounded-SAM reducing the number of proposals, our approach remains slower compared to end-to-end detectors. This is primarily due to the process involved in obtaining proposal embeddings. The inclusion of adapters does not impede the overall framework. Future research could focus on developing a more efficient method for robotic applications. Please see the appendix for segmentation runtime.
4.4 Ablation Study
Grounded SAM (GS) vs SAM On the seven BOP datasets containing images of numerous cluttered scenes, we compare the object proposals from GS and SAM as detailed in Table 5. The results indicate that GS yields more precise bounding boxes and masks compared to using SAM alone. Furthermore, GS enhances efficiency by eliminating false object proposals, thereby reducing runtime.
FFA vs $cls$ token To compare these two types of embedding generation, we present the results in Table 5. Our weight adapter enhances both two types of embeddings. Despite close segmentation results, FFA produces embeddings that possess greater adaptive potential, demonstrated by higher AP scores after adaption via our weight adapter.
Table 5: Comparison of segmentation performance across various proposal models and embedding generation options. The first row is from CNOS [64].
Proposal Embedding BOP Datasets
LM-O T-LESS TUD-L IC-BIN ITODD HB YCB-V Mean
SAM cls token 39.6 39.7 39.1 28.4 28.2 48.0 59.5 40.4
SAM FFA 40.0 40.0 40.8 28.0 25.8 46.0 54.4 39.3
GS cls token 41.7 41.7 50.8 31.5 30 58 63 45.2
GS FFA 42.9 43.0 52.0 30.5 28.8 56.6 59.7 44.8
GS cls token + WA 42.5 48.2 50.7 31.3 32.0 60.5 62.8 46.9
GS FFA +WA 43.2 49.2 53.5 30.5 29.4 62.4 64.3 47.5
Table 6: Detection results using different ViT backbones of Dinov2. “reg” indicates DINOv2 with registers. The results are based on all testing images of the high-resolution dataset. WA Diff indicates the improvement attributed to the Weight Adapter.
Dinov2 backbone AP AP50 AP75 WA Diff
ViT-S/14 47.5 56.4 51.7 + 0.9 AP
ViT-S/14 + WA 48.4 57.5 52.9
ViT-B/14 54.7 65.2 60.1 + 3.2 AP
ViT-B/14 + WA 57.9 69.0 63.6
ViT-L/14 56.8 67.7 62.3 + 3.6 AP
ViT-L/14 + WA 60.4 71.8 66.4
ViT-L/14 reg 59.3 71.1 65.1 + 4.6 AP
ViT-L/14 reg +WA 63.9 76.6 70.6
Dinov2 backbones with adapter Our Weight Adapter is compatible with various backbones of Dinov2. Notably, more powerful backbones, which offer a more effective feature space, enable our adapter to deliver greater improvements. Details of this comparison are provided in Table 6. Please refer to the Appendix for additional ablation studies.
5 Discussions
In this study, we introduce NIDS-Net, a framework designed for novel instance detection and segmentation. We utilize Grounding DINO and SAM to generate precise foreground object proposals, as opposed to conventional naive region proposals. We also present the Weight Adapter, which refines features from a pre-trained vision model effectively and mitigates the risk of overfitting. With the adapter, template and proposal embeddings of different instances are separated to facilitate the subsequent matching. Our method surpasses other approaches significantly in detection performance and also excels in segmentation compared to existing methods.
However, there are limitations: given that our approach incorporates multiple pre-trained models, it requires a longer running time and greater computational resources compared to end-to-end detectors. Additionally, Grounding DINO may produce large bounding boxes that encompass multiple objects, leading to under-segmentation and detection failures. Moreover, our method tends to overlook heavily occluded objects with low confidence scores.
In this work, instances are represented by $K$ template embeddings. For future research, we propose that each instance should be characterized by a single, distinctive embedding that acts as its identifier to enable one-shot detection. This approach will allow the detector to identify and locate a target instance within a query image using just one template image. Additionally, developing an efficient method for this process will be crucial for its application in robotics.
Acknowledgments
This work was supported in part by the DARPA Perceptually-enabled Task Guidance (PTG) Program under contract number HR00112220005.

### v2

@@
 Novel Instance Detection and Segmentation (NIDS) aims at detecting and segmenting novel object instances given a few examples of each instance.
-We propose a unified framework (NIDS-Net) comprising object proposal generation, embedding creation for both instance templates and proposal regions, and embedding matching for instance label assignment.
-Leveraging recent advancements in large vision methods, we utilize the Grounding DINO and Segment Anything Model (SAM) to obtain object proposals with accurate bounding boxes and masks.
+We propose a unified, simple yet effective framework (NIDS-Net) comprising object proposal generation, embedding creation for both instance templates and proposal regions, and embedding matching for instance label assignment.
+Leveraging recent advancements in large vision methods, we utilize Grounding DINO and Segment Anything Model (SAM) to obtain object proposals with accurate bounding boxes and masks.
 Central to our approach is the generation of high-quality instance embeddings.
 We utilize foreground feature averages of patch embeddings from the DINOv2 ViT backbone, followed by refinement through a weight adapter mechanism that we introduce.
-We show experimentally that our weight adapter can adjust the embeddings locally within their feature space and effectively limit overfitting.
+We show experimentally that our weight adapter can adjust the embeddings locally within their feature space and effectively limit overfitting in the few-shot setting.
 This methodology enables a straightforward matching strategy, resulting in significant performance gains.
 Our framework surpasses current state-of-the-art methods, demonstrating notable improvements of 22.3, 46.2, 10.3, and 24.0 in average precision (AP) across four detection datasets.
-In instance segmentation tasks on seven core datasets of the BOP challenge, our method outperforms the top RGB methods by 3.6 AP and remains competitive with the best RGB-D method [^1] Code is available at: https://github.com/YoungSean/NIDS-Net.
+In instance segmentation tasks on seven core datasets of the BOP challenge, our method is around 4.5 times faster than the leading published RGB method and surpasses it by 3.6 AP.
+NIDS-Net is about 5.7 times faster than the top RGB-D method while maintaining competitive performance [^1] Project Page: https://irvlutd.github.io/NIDSNet/.
 1 Introduction
+Figure 1: We leverage pre-trained vision models for object proposal generation and feature extraction, and introduce a weight adapter to improve pre-trained feature embeddings for novel object instance detection and segmentation.
 Novel Instance Detection and Segmentation (NIDS) is a crucial task in computer vision aimed at identifying and locating unseen instances in images or videos, given a few examples of each instance.
-Of particular interest in this work, are applications of NIDS in robotics, e.g., a robot might need to fetch a specific, novel object instance from a cluttered desk given only a small number of multi-view template images of the novel object.
-Ideally, detectors for this task should provide precise bounding boxes and masks of the target instances in a cluttered query image, but standard detection systems, trained on specific classes, often struggle to recognize new objects.
-The current paradigm for NIDS typically encompasses the following steps: (1) generating proposals from a query image, (2) obtaining embeddings of the proposals and the target instance templates, and (3) matching the embeddings of proposals with those of the templates for identification.
+It also has wide applications in robotics.
+For example, suppose a robot needs to grasp a specific, novel object instance from a cluttered desk given only a small number of multi-view template images of the object.
+Ideally, detectors for this task should provide precise bounding boxes and masks of the target instances given a query image.
+However, standard detection methods trained on specific classes often struggle to recognize new objects.
+The current paradigm for NIDS typically encompasses the following steps: (1) generating proposals from a query image, (2) obtaining embeddings of the proposals and the object templates, and (3) matching the embeddings of proposals with those of the templates for identification.
 Recent work [49, 81, 66, 53] has utilized various open-world detectors to obtain object proposals.
 However, the use of models such as the Segment Anything Model (SAM) [47] or FastSAM [103] often results in the generation of region-based proposals rather than actual object proposals.
-This misidentification of regions as objects can to lead to object identification issues.
+This misidentification of regions as objects can lead to object identification issues.
 For example, a single object may be divided into multiple proposals or some background regions may be mistaken for foreground objects.
 These false alarms not only slow down the whole detection progress but also spoil the final results.
-To represent proposals and templates, Shen et al.
-[79], Nguyen et al.
-[64], and Lin et al.
-[49] adopt the $cls$ token of DINOv2  [66, 20] while Li et al.
-[45] employ a 3D voxel representation.
+To embed proposals and templates, some existing works, e.g., [81, 66, 53], adopt the $cls$ token of DINOv2  [68, 20] while others, e.g., Li et al.
+[49], employ a 3D voxel representation.
 Ideally, for each specific unseen instance, the template embeddings from different views should be similar to each other but markedly different from the embeddings of other instances.
-Despite these efforts, the challenge persists that the embeddings of different instances remain excessively similar.
-The development of robust feature embeddings could provide a solution to this issue.
-To address the limitations of current methods, we propose a framework for NIDS.
+Despite these efforts, the embeddings of different instances may remain similar, and so the challenges persist.
+To address the limitations of the existing methods, we propose a framework for NIDS, as depicted in Fig.
+1.
 Initially, we utilize Grounding DINO [56] with the text prompt “objects” on a cluttered query image to obtain bounding boxes for foreground objects.
@@
 Subsequently, we employ SAM to create masks within these bounding boxes, generating precise object proposals comprising both bounding boxes and masks.
-For instance embeddings, we first extract average foreground features from the patch embeddings of the DINOv2 ViT backbone.
+For instance embeddings, we first extract average foreground features [48] from the patch embeddings of the DINOv2 ViT backbone.
 We then apply adapters to enhance instance embeddings by clustering similar instances and distancing different ones.
-The adapters are trained with instance template embeddings via InfoNCE loss [65, 15].
+The adapters are trained using the template images only, where InfoNCE loss [67, 15] is applied to the feature embeddings after the adapters.
 To refine these embeddings, the CLIP-Adapter [29] introduces a residual vector added to the original embedding, which risks overfitting with a few training examples.
 This addition can spoil and destabilize the embeddings of background objects, disrupting the entire feature space.
-To mitigate this issue, we introduce a weight adapter that modifies the original embeddings by applying learned weights.
+To mitigate this issue, we, instead, introduce a weight adapter that modifies the original embeddings by applying learned weights.
 Finally, we employ a straightforward matching approach [62], such as stable matching or using $argmax$, to assign instance labels to these proposals.
-Our approach has been validated on four detection datasets and seven segmentation datasets.
+Our approach has been validated on four detection datasets and seven segmentation datasets for NIDS.
 The framework significantly surpasses existing state-of-the-art methods: it demonstrates substantial increases in average precision (AP), with gains of 22.3, 46.2, 10.3, and 24.0 across these detection datasets.
-Moreover, in instance segmentation tasks on the seven core datasets of the BOP challenge [82], our method outperforms the best RGB and RGB-D methods.
+Moreover, in instance segmentation tasks on the seven core datasets of the BOP challenge [84], our method outperforms the best published RGB and RGB-D methods.
 Our contributions are summarized as follows.
-  • We propose NIDS-Net, a unified framework for novel instance detection and segmentation that surpasses existing methods across four detection datasets and seven segmentation datasets of BOP challenge.
-  • We leverage the objectness of open-world detector Grounding DINO and region recognition of Segmentation Anything Model (SAM) to obtain object proposals.
+  • We propose NIDS-Net, a unified framework for novel instance detection and segmentation that surpasses existing published methods across four detection datasets and seven segmentation datasets of BOP challenge.
+  • We leverage the objectness of the open-world detector Grounding DINO and region recognition of Segmentation Anything Model (SAM) to obtain object proposals.
   • We introduce the Weight Adapter, a tool designed to refine embeddings within their feature space while discouraging overfitting.
@@
 Additionally, [3, 4, 96] address depth prediction, and [72, 73, 6, 87] tackle large-scale language modeling.
-Object detection is addressed by [30, 77, 75, 40, 41, 31, 52], while image segmentation is the focus of [33, 86, 42].
+Object detection is addressed by [79, 77, 43, 44, 31, 56], while image segmentation is the focus of [33, 89, 46].
 Works such as [94, 52, 14, 76] explore text-to-3D generation, and [82, 1] address text-to-video synthesis.
 Finally, 6DoF pose estimation and tracking are tackled by [90, 21, 51, 88].
+DINOv2 [68, 20] offers robust features to represent unseen instances.
 These foundational models offer diverse capabilities, with the challenge lying in effectively leveraging their wealth of knowledge for specific use cases, such as novel instance detection.
-DINOv2 [66, 20] offers robust features to represent unseen instances.
+In our work, we leverage pre-trained vision models for NIDS.
 Open-World Object Detection (OWOD).
-OWOD tackles real-world scenarios where object detection algorithms are tasked not only with identifying known and labeled objects but also with effectively handling novel and unknown objects absent during inference.
+OWOD tackles real-world scenarios where object detection algorithms are tasked not only with identifying known and labeled objects but also with effectively handling novel and unknown objects during inference.
 This sets OWOD apart from traditional object detection, which confines itself to detecting only predetermined object classes.
@@
 Recent approaches such as OWL-ViT [65], DITO [43], RO-ViT [44], YOLO-World [18], Sambor [31], and Grounding DINO [56] offer diverse solutions to address OWOD challenges.
-OWL-ViT leverages a standard Vision Transformer [24] architecture with minimal modifications, incorporating contrastive image-text pre-training and end-to-end detection fine-tuning.
+For example, OWL-ViT leverages a standard Vision Transformer [24] architecture with minimal modifications, incorporating contrastive image-text pre-training and end-to-end detection fine-tuning.
 DITO introduces an open-vocabulary detection approach based on detection-oriented image-text pretraining, replacing the classification architecture with a detector architecture during pretraining to better serve region-level recognition needs.
-RO-ViT proposes cropping and resizing regions of positional embeddings (PE) at pretraining, aligning better with region-level detection fine-tuning.
-YOLO-World enhances YOLO [75] with open vocabulary detection capabilities through vision-language modeling and large-scale pretraining.
-Grounding DINO [52] utilizes features from DINO [10] to ground text queries, while Sambor extends SAM [42] to detect arbitrary objects based on human inputs such as category names or reference expressions.
+Grounding DINO [56] utilizes features from DINO [10] to ground text queries.
 We employ Grounding DINO due to its strong performance.
+Similar to our work, Fang et al.
+[28] obtain object bounding boxes using the text prompt “a rigid object” with Grounding DINO.
 Few-Shot Object Detection (FSOD).
 FSOD involves detecting occurrences of a novel class, unseen during training, given only a few template examples.
-De-ViT [100] introduces a region-propagation mechanism for localization, transforming propagated region masks into bounding boxes using a learnable spatial integral layer.
+For example, De-ViT [102] introduces a region-propagation mechanism for localization, transforming propagated region masks into bounding boxes using a learnable spatial integral layer.
 FS-DETR [7], the first few-shot detection transformer, extends DETR [9] to process arbitrary numbers of novel objects concurrently while supporting various examples from each class.
@@
 Additionally,  [60] generates specific and representative prototypes for each query image, considering both support and query image features.
+Our method is a FSOD method based on region proposal generation and matching.
+Compared to previous FSOD methods, it is simple yet effective.
 Instance Detection.
@@
 We generate 2D robust instance embeddings from templates with DINOv2 [68, 20].
-Unseen object instance segmentation (UOIS).
-UOIS is a task designed to segment unseen objects in cluttered scenes.
-UOIS generates object masks without providing labels.
-UOIS-Net [89] utilizes a two-stage framework, initially generating rough masks using depth information, and then refining them into precise instance masks with RGB data.
-Both UCN [88] and MSMFormer [57, 58] segment objects by clustering pixel embeddings.
-Although these methods can generate object proposals, they primarily focus on tabletop settings.
-In contrast, our approach is designed to handle RGB query images in various environments.
-Adapter.
+Adapters for Pre-trained Models.
 The application of adapters atop large pre-trained models has emerged as a prominent strategy, yielding significant improvements across various tasks.
@@
 The adapter assigns weights to the embeddings from frozen backbones.
+SENets [38] employ the Squeeze-and-Excitation (SE) blocks to distribute weights.
+In contrast, our adapter is an independent model with a structural difference from the SE block.
+For instance, the SE block lacks a ReLU layer prior to the sigmoid function.
 3 Method
@@
 We will assume that each of the $N$ target instances is represented by $K$ template images $I_{T}\in\mathbb{R}^{K\times 3\times W\times H}$ and their corresponding segmentation masks.
-During inference, the output of each query image $I_{Q}\in\mathbb{R}^{3\times W\times H}$ provides these instances’ bounding boxes for detection and instance masks for segmentation.
+During inference, the output of each query image $I_{Q}\in\mathbb{R}^{3\times W\times H}$ provides the bounding boxes for detection and instance masks for segmentation of these instances.
 3.1 Instance Embedding Generation Stage
@@
 Each of the $N$ instances has $K$ multi-view template images and their ground truth segmentation masks from which we derive template embeddings $E_{T}\in\mathbb{R}^{N\times K\times C}$, where $C$ is the dimension of the embeddings.
-Given an image and its corresponding mask, we initially extract patch embeddings using the ViT backbone of DINOv2, and subsequently obtain foreground features as specified by the mask.
+Given an image and its corresponding mask, we initially extract patch embeddings using the ViT backbone of DINOv2 [68], and subsequently obtain foreground features as specified by the mask.
 We then perform average pooling on these features.
 This process, termed Foreground Feature Averaging (FFA), is proposed by [48] to assess object similarity.
-We employ FFA to generate all initial instance embeddings.
+We employ FFA to generate all initial instance embeddings using the object templates.
 3.2 Object Proposal Stage
@@
 However, some embeddings from distinct instances may cluster together.
-For exxample, background object embeddings might be similar to targets.
+For example, background object embeddings might be similar to targets.
 Some target instances may resemble each other.
 To address this issue, we employ learnable adapters to refine the embeddings.
-We train these adapters using InfoNCE loss [65, 15], aiming to bring the embeddings of the same instance closer together while separating the embeddings of different instances.
+We train these adapters using the InfoNCE loss [67, 15], aiming to bring the embeddings of the same instance closer together while separating the embeddings of different instances.
+This training only uses the template images.
 CLIP-Adapter (CA).
@@
 To mitigate this issue, CLIP-Adapter integrates residual connections that dynamically merge the newly adapted knowledge with the foundational knowledge from the original CLIP backbone.
-Given an instance embedding $f$ and a $MLP$ (Multi-Layer Perceptron), the CLIP-Adapter modifies the embedding according to the formula $f_{c}=\alpha\times MLP(f)+(1-\alpha)\times f$.
+Given an instance embedding $f$ and an $MLP$ (Multi-Layer Perceptron), the CLIP-Adapter modifies the embedding according to the formula $f_{c}=\alpha\times MLP(f)+(1-\alpha)\times f$.
 In this equation, $f_{c}$ represents the adapted embedding, which is a linear combination of the transformed embedding $MLP(f)$ and the original embedding $f$.
@@
 Given that cosine similarity is utilized to compare embeddings in the subsequent phase, our adapter alters the interrelationships among the values within each embedding.
-Further details of adapters are presented in Fig.
-2.
-Qualitative results can be found in the appendix.
+Further adapter details are presented in Fig.
+3.
 Figure 3: Comparison of two adapters. (left) CLIP-Adapter. (right) Our weight adapter.
@@
 It yields a matrix of template scores with dimensions $Q\times N\times K$, as illustrated in Fig.
-1.
+2.
 For each instance, we aggregate all $K$ template scores to derive a matrix of instance scores with dimensions $Q\times N$.
@@
 To improve segmentation performance, we incorporate an additional appearance matching score, $s_{\text{appe}}$, as proposed by SAM-6D [53], into the instance scores.
+The final instance scores are computed as the average of $s_{\text{appe}}$ and the initial instance scores.
 $s_{\text{appe}}$ is used to identify objects that are semantically similar yet differ in appearance.
@@
 $s_{\text{appe}}$ is derived from the patch embeddings of a proposal image $I$ and $T_{\text{best}}$.
-It quantifies the similarity between the query image and the template in terms of their local features as follows:
-$s_{\text{appe}}=\frac{1}{N_{\text{patch}}^{I}}\sum_{j=1}^{N_{\text{patch}}^{I}}\max_{i=1,\ldots,N_{\text{patch}}^{T_{\text{best}}}}\left(CosineSimilarity(f_{I,j}^{\text{patch}},f_{T_{\text{best}},i}^{\text{patch}})\right)$
+It quantifies the similarity between the query image and the best template in terms of their local features ($f_{I,j}^{\text{patch}}$, $f_{T_{\text{best}},i}^{\text{patch}}$) as follows:
+$s_{\text{appe}}=\frac{1}{N_{I}^{\text{patch}}}\sum_{j=1}^{N_{I}^{\text{patch}}}\max_{i=1,\ldots,N_{T_{\text{best}}}^{\text{patch}}}\left(\frac{f_{I,j}^{\text{patch}}\cdot f_{T_{\text{best}},i}^{\text{patch}}}{\|f_{I,j}^{\text{patch}}\|_{2}\|f_{T_{\text{best}},i}^{\text{patch}}\|_{2}}\right),$ (3)
+where $N^{\text{patch}}$ is the number of patch embeddings.
 Assuming that objects are unique within a query image [81], we employ the stable matching algorithm [62] on instance scores to assign a unique instance ID to each proposal.
@@
 Each is defined by $\{b_{i},M_{i},o_{i},s_{i}\}$.
-Here, $b_{i}$ is the bounding box of an instance.
-$M_{i}$ is the modal mask, which covers the visible object surface [82].
-$o_{i}$ is the instance ID.
-$s_{i}$ is the confidence score.
+Here, $b_{i}$ is the bounding box of an instance $M_{i}$ is the modal mask, which covers the visible object surface [84], $o_{i}$ is the instance ID, and $s_{i}$ is the confidence score.
 A confidence score threshold $\delta$ can be used to remove incorrect proposals.
-4 Experiments
-We employ the Grounding DINO model with a Swin-T backbone [52] and a text prompt “objects”, and the default ViT-H SAM [43] to generate object proposals.
-According to the proposals, we obtain the regions of interest (RoIs) and resize them to $448\times 448$ or $224\times 224$ resolutions.
-For object detection, instance embeddings are produced using the DINOv2’s ViT-L model with registers [20], employing stable matching due to the uniqueness of instances in these datasets.
-For image segmentation, following SAM-6D [49], we utilize the ViT-L model of DINOv2 [66].
-In scenarios with identical instances, such as the cluttered scenes in BOP datasets, we apply the $argmax$ function for matching.
-The experiments are conducted on an RTX A5000 GPU.
-For each dataset, we train both adapters on template embeddings using InfoNCE loss [65, 15] with temperature 0.05.
-With two linear layers, Both adapters initially decrease the embedding dimension from $C$ to $C/4$, subsequently restoring it to $C$.
-The Weight Adapter is trained with Adam optimizer at a learning rate of 1e-3 and a batch size of 1024, while the CLIP-Adapter is trained at a learning rate of 1e-4.
-Dropout is incorporated in the CLIP-Adapter to reduce overfitting.
-For additional details, please see the Appendix.
-4.1 Detection Datasets
-We utilize recently developed instance detection datasets and their associated baselines [79, 45].
-The High-resolution and RoboTools datasets employ real template images, while the LM-O and YCB-V datasets utilize synthetic images.
-High-resolution dataset Shen et al.
-[79] design a high-resolution dataset for instance detection.
-This dataset comprises 100 distinct object instances, each represented by 24 photos taken from multiple viewpoints, with each photo having a resolution of $3072\times 3072$ pixels.
-These instances are integrated into 14 different indoor scenes for testing, captured in even higher resolution ($6144\times 8192$ pixels).
-The testing images are further classified as easy or hard based on the level of scene clutter and how much the instances are occluded.
-Easy tags are assigned when objects are sparsely placed, while hard tags are used for more cluttered setups.
-Baselines According to Shen et al.
-[79], two types of baselines are set for the high-resolution dataset: Cut-Paste-Learn and a non-learned method.
-Cut-Paste-Learn generates synthetic training images with 2D-box annotations, putting foreground instances in various sizes and aspect ratios onto different backgrounds.
-This enables training detectors by viewing each instance as a class.
-They train five detectors: FasterRCNN [76], RetinaNet [51], CenterNet [102], FCOS [83], and the transformer-based DINO [96].
-For the non-learned method, SAM is initially used to generate proposals, followed by employing DINO [11] and DINOv2 [66] to generate features for both proposal and template images, ultimately performing proposal matching and selection.
-Synthetic-Real test sets Li et al.
-[45] employ two benchmarks for evaluation.
-LineMod-Occlusion (LM-O) [5] includes 8 texture-less objects and 1,514 bounding boxes, primarily challenging due to significant occlusion.
-The YCB-Video (YCB-V) [8] features 21 objects and 4,125 bounding boxes, with pose variation as the main challenge.
-Since these datasets have real testing images without reference videos, Li et al.
-[45] generate synthetic videos using CAD models.
-We sample 16 synthetic template images per object from these videos.
-RoboTools benchmark Li et al.
-[45] introduce RoboTools, a challenging real-world benchmark for robotics.
-It features 20 distinct instances, 9,109 annotations, and 24 complex scenarios.
-Compared to previous benchmarks [5, 8], RoboTools offers greater difficulty due to its cluttered backgrounds and significant pose variations.
-Additionally, the benchmark’s reference videos are made from real images and incorporate realistic lighting conditions, including shadows.
-25 real template images of each instance are sampled from reference videos.
-Baselines For cluttered scenes from the datasets RoboTools, LM-O and YCB-V, Li et al.
-[45] have developed several 2D baselines: OLNDINO, OLNCLIP, and OLNCorr.
-Initially, they generate open-world 2D proposals using their detection module [39].
-For matching, different methods are employed to select the proposal with the highest score.
-In OLNDINO and OLNCLIP, they use robust features from pre-trained backbones [11, 73] and cosine similarity for matching.
-For OLNCorr, a matching head is designed based on correlation [53].
-They also employ the class-level one-shot detectors OS2D [67] and BHRL [94].
-To address the limitations of traditional 2D methods with pose variations and occlusions, Li et al.
-[45] introduced VoxDet which is based on a 3D voxel representation.
-To train these methods, Li et al.
-[45] developed the open-world instance detection (OWID-10k) dataset as a synthetic instance detection training set.
-It features 9,901 objects from ShapeNet [12] and ABO [19], each rendered into a 40-frame, 360-degree video using Blenderproc [22].
-Evaluation metrics For detection, we evaluate our method with Average Precision (AP).
-AP is computed by averaging precision scores at various Intersection over Union (IoU) thresholds, specifically from 0.5 to 0.95, in increments of 0.05.
-This metric is the primary evaluation standard in the COCO Object Detection dataset [50].
-Additionally, AP50 and AP75 are variations of this metric, where precision is averaged across all instances at IoU thresholds of 0.5 and 0.75, respectively.
-4.2 Segmentation Datasets
-BOP challenge We test our method using seven datasets from the BOP challenge [82]: LineMod Occlusion (LM-O) [5], T-LESS [34], TUD-L [35], IC-BIN [25], ITODD [26], HomebrewedDB (HB) [38], and YCB-Video [8].
-The objects from these datasets vary widely.
-We use 42 template rendering images from CNOS [64], which are generated via BlenderProc [22].
-Baselines We compare our method with ZeroPose [13], CNOS [64], and SAM-6D [49].
-These methods utilize proposals from SAM [43] or FastSAM [101] and use the $cls$ token from DINOv2 as the instance embedding for matching.
-In addition to adjusting the SAM hyperparameters to generate more proposals, SAM-6D enhances its performance by incorporating an appearance score, $s_{\text{appe}}$.
-Moreover, SAM-6D employs a Geometric Matching Score, $s_{\text{geo}}$, which considers the shapes and sizes of instances during matching by utilizing depth information.
-Evaluation metrics For instance segmentation task, we evaluate our method using the Average Precision (AP) metrics.
-The AP is computed by averaging the precision scores at various IoU thresholds, from 0.50 to 0.95, increasing by 0.05 at each step.
-Figure 3: Visual examples of our results using the Weight Adapter on the high-resolution and RoboTools datasets.
-Our approach detects specific object instances in cluttered scenes according to their real template images.
 Table 1: Detection results on the high-resolution dataset.
@@
 NIDS-Net + WA (Ours) 63.9 43.4 72.7 18.1 62.5 84.0 76.6 70.6 6.78
+4 Experiments
+We employ the Grounding DINO model with a Swin-T backbone [56] and a text prompt “objects”, and the default ViT-H SAM [47] to generate object proposals.
+According to the proposals, we obtain the regions of interest (RoIs) and resize them to $448\times 448$ or $224\times 224$ resolutions.
+For object detection, instance embeddings are produced using the DINOv2’s ViT-L model with registers [20], employing stable matching due to the uniqueness of instances in these datasets.
+For image segmentation, following SAM-6D [53], we utilize the ViT-L model of DINOv2 [68].
+In scenarios with identical instances, such as the cluttered scenes in BOP datasets, we apply the $argmax$ function for matching.
+The experiments are conducted on an RTX A5000 GPU.
+For each dataset, we train both adapters on template embeddings using InfoNCE loss [67, 15] with temperature 0.05.
+With two linear layers, both adapters initially decrease the embedding dimension from $C$ to $C/4$, subsequently restoring it to $C$.
+The Weight Adapter is trained with the Adam optimizer [45] at a learning rate of 1e-3 and a batch size of 1024, while the CLIP-Adapter is trained at a learning rate of 1e-4.
+Dropout [34] is incorporated in the CLIP-Adapter to reduce overfitting.
+For additional details, please see the appendix.
+4.1 Detection Datasets
+We utilize recently developed instance detection datasets and their associated baselines [81, 49].
+The High-resolution and RoboTools datasets employ real template images, while the LM-O and YCB-V datasets utilize synthetic images.
+High-resolution Dataset.
+Shen et al.
+[81] design a high-resolution dataset for instance detection.
+This dataset comprises 100 distinct object instances, each represented by 24 photos taken from multiple viewpoints, with each photo having a resolution of $3072\times 3072$ pixels.
+These instances are integrated into 14 different indoor scenes for testing, captured in even higher resolution ($6144\times 8192$ pixels).
+The testing images are further classified as easy or hard based on the level of scene clutter and how much the instances are occluded.
+Easy tags are assigned when objects are sparsely placed, while hard tags are used for more cluttered setups.
+According to Shen et al.
+[81], two types of baselines are set for the high-resolution dataset: Cut-Paste-Learn and a non-learned method.
+Cut-Paste-Learn generates synthetic training images with 2D-box annotations, putting foreground instances in various sizes and aspect ratios onto different backgrounds.
+This enables training detectors by viewing each instance as a class.
+They train five detectors: FasterRCNN [78], RetinaNet [55], CenterNet [104], FCOS [85], and the transformer-based DINO [98].
+For the non-learned method, SAM is initially used to generate proposals, followed by employing DINO [11] and DINOv2 [68] to generate features for both proposal and template images, ultimately performing proposal matching and selection.
+Synthetic-Real Test Sets.
+Li et al.
+[49] employ two benchmarks for evaluation.
+LineMod-Occlusion (LM-O) [5] includes 8 texture-less objects and 1,514 bounding boxes, primarily challenging due to significant occlusion.
+The YCB-Video (YCB-V) [8] features 21 objects and 4,125 bounding boxes, with pose variation as the main challenge.
+Since these datasets have real testing images without reference videos, Li et al.
+[49] generate synthetic videos using CAD models.
+We sample 16 synthetic template images per object from these videos.
+RoboTools Benchmark.
+Li et al.
+[49] introduce RoboTools, a challenging real-world benchmark for robotics.
+It features 20 distinct instances, 9,109 annotations, and 24 complex scenarios.
+Compared to previous benchmarks [5, 8], RoboTools offers greater difficulty due to its cluttered backgrounds and significant pose variations.
+Additionally, the benchmark’s reference videos are made from real images and incorporate realistic lighting conditions, including shadows.
+25 real template images of each instance are sampled from reference videos.
+For cluttered scenes from the datasets RoboTools, LM-O and YCB-V, Li et al.
+[49] have developed several 2D baselines: OLNDINO, OLNCLIP, and OLNCorr.
+Initially, they generate open-world 2D proposals using their detection module [42].
+For matching, different methods are employed to select the proposal with the highest score.
+In OLNDINO and OLNCLIP, they use robust features from pre-trained backbones [11, 75] and cosine similarity for matching.
+For OLNCorr, a matching head is designed based on correlation [57].
+They also employ the class-level one-shot detectors OS2D [69] and BHRL [95].
+To address the limitations of traditional 2D methods with pose variations and occlusions, Li et al.
+[49] introduced VoxDet which is based on a 3D voxel representation.
+To train these methods, Li et al.
+[49] developed the open-world instance detection (OWID-10k) dataset as a synthetic instance detection training set.
+It features 9,901 objects from ShapeNet [12] and ABO [19], each rendered into a 40-frame, 360-degree video using Blenderproc [22].
+Evaluation Metrics.
+For detection, we evaluate our method with Average Precision (AP).
+AP is computed by averaging precision scores at various Intersection over Union (IoU) thresholds, specifically from 0.5 to 0.95, in increments of 0.05.
+This metric is the primary evaluation standard in the COCO Object Detection dataset [54].
+Additionally, AP50 and AP75 are variations of this metric, where precision is averaged across all instances at IoU thresholds of 0.5 and 0.75, respectively.
+4.2 Segmentation Datasets
+The BOP Challenge.
+We test our method using seven datasets from the BOP challenge [84]: LineMod Occlusion (LM-O) [5], T-LESS [35], TUD-L [36], IC-BIN [25], ITODD [26], HomebrewedDB (HB) [40], and YCB-Video [8].
+The objects from these datasets vary widely.
+We use 42 template rendering images from CNOS [66], which are generated via BlenderProc [22].
+Baselines.
+We compare our method with ZeroPose [13], CNOS [66], and SAM-6D [53].
+These methods utilize proposals from SAM [47] or FastSAM [103] and use the $cls$ token from DINOv2 as the instance embedding for matching.
+In addition to adjusting the SAM hyperparameters to generate more proposals, SAM-6D enhances its performance by incorporating an appearance score, $s_{\text{appe}}$.
+Moreover, SAM-6D employs a Geometric Matching Score, $s_{\text{geo}}$, which considers the shapes and sizes of instances during matching by utilizing depth information.
+Evaluation Metrics.
+For instance segmentation task, we evaluate our method using the Average Precision (AP) metrics.
+The AP is computed by averaging the precision scores at various IoU thresholds, from 0.50 to 0.95, increasing by 0.05 at each step.
+Figure 4: Visual examples of our results using the Weight Adapter on the high-resolution and RoboTools datasets.
+Our approach detects specific object instances in cluttered scenes according to their real template images.
 Table 2: Detection performance on the fully real dataset, RoboTools.
@@
 †means the model is trained on OWID and real images.
-Our running time of YCB-V is based on the scene 48.
 ()($\cdot$)( ⋅ ) represents the number of instances.
@@
 ()($\cdot$)( ⋅ ) includes the number of instances.
+SAM6D (RGBD) exclusively utilizes RGB-D images, whereas other models employ only RGB images.
 Method BOP Datasets
@@
 NIDS-Net + WA(Ours) 43.2 49.2 53.5 30.5 29.4 62.4 64.3 47.5
-NIDS-Net + WA (Ours) + $s_{\text{appe}}$ 43.9 49.6 55.6 32.8 31.5 62.0 65.0 48.6
+NIDS-Net + WA + $s_{\text{appe}}$ (Ours) 43.9 49.6 55.6 32.8 31.5 62.0 65.0 48.6
+Table 5: Comparison of segmentation performance across various proposal models and embedding generation options. The first row is from CNOS [66].
+Proposal Embedding BOP Datasets
+LM-O T-LESS TUD-L IC-BIN ITODD HB YCB-V Mean
+SAM cls token 39.6 39.7 39.1 28.4 28.2 48.0 59.5 40.4
+SAM FFA 40.0 40.0 40.8 28.0 25.8 46.0 54.4 39.3
+GS cls token 41.7 41.7 50.8 31.5 30 58 63 45.2
+GS FFA 42.9 43.0 52.0 30.5 28.8 56.6 59.7 44.8
+GS cls token + WA 42.5 48.2 50.7 31.3 32.0 60.5 62.8 46.9
+GS FFA +WA 43.2 49.2 53.5 30.5 29.4 62.4 64.3 47.5
+Figure 5: Comparison of segmentation results using CNOS, SAM6D, and our NIDS-Net on the YCB-V dataset.
+CNOS and SAM6D may misclassify some background regions or object parts as objects due to their reliance on SAM for proposal generation.
+NIDS-Net addresses this limitation with Grounded-SAM.
 4.3 Benchmarking Results
-Detection The high-resolution dataset results are presented in Table 1.
+Detection.
+The high-resolution dataset results are presented in Table 1.
 Our models dramatically outperform existing techniques, achieving the highest AP scores across all categories.
@@
 The improved performance on YCB-V, which contains more instances, indicates that our adapter functions more effectively with increased instance variety.
-Segmentation CNOS [64] and SAM-6D [49] initially utilize SAM or FastSAM to obtain proposals and derive instance embeddings through the $cls$ token of Dinov2.
+Segmentation.
+CNOS [66] and SAM-6D [53] initially utilize SAM or FastSAM to obtain proposals and derive instance embeddings through the $cls$ token of Dinov2.
 In contrast, we generate proposals using Grounded-SAM and acquire embeddings through the FFA pipeline.
@@
 Overall, our method surpasses the state-of-the-art SAM-6D and achieves superior results compared to RGBD results.
-Qualitative result We present detection results for both simple and complex scenes in Fig. 3. Additional detection and segmentation examples are available in the Appendix.
-Runtime The tables of detection results include the runtime of our method.
-Despite Grounded-SAM reducing the number of proposals, our approach remains slower compared to end-to-end detectors.
-This is primarily due to the process involved in obtaining proposal embeddings.
-The inclusion of adapters does not impede the overall framework.
-Future research could focus on developing a more efficient method for robotic applications.
+Qualitative Result.
+We present detection results for both simple and complex scenes in Fig.
+4.
+Fig.
+5 displays a visual comparison of the segmentation results.
+Additional examples are available in the appendix.
+Runtime.
+The tables of detection results include the runtime of our method.
+Grounded-SAM substantially decreases the number of proposals compared to SAM, thereby accelerating the entire process.
 Please see the appendix for segmentation runtime.
 4.4 Ablation Study
-Grounded SAM (GS) vs SAM On the seven BOP datasets containing images of numerous cluttered scenes, we compare the object proposals from GS and SAM as detailed in Table 5.
+Grounded SAM (GS) vs SAM.
+On the seven BOP datasets containing images of numerous cluttered scenes, we compare the object proposals from GS and SAM as detailed in Table 5.
 The results indicate that GS yields more precise bounding boxes and masks compared to using SAM alone.
 Furthermore, GS enhances efficiency by eliminating false object proposals, thereby reducing runtime.
-FFA vs $cls$ token To compare these two types of embedding generation, we present the results in Table 5.
+FFA vs $cls$ token.
+To compare these two types of embedding generation, we present the results in Table 5.
 Our weight adapter enhances both two types of embeddings.
 Despite close segmentation results, FFA produces embeddings that possess greater adaptive potential, demonstrated by higher AP scores after adaption via our weight adapter.
-Table 5: Comparison of segmentation performance across various proposal models and embedding generation options. The first row is from CNOS [64].
-Proposal Embedding BOP Datasets
-LM-O T-LESS TUD-L IC-BIN ITODD HB YCB-V Mean
-SAM cls token 39.6 39.7 39.1 28.4 28.2 48.0 59.5 40.4
-SAM FFA 40.0 40.0 40.8 28.0 25.8 46.0 54.4 39.3
-GS cls token 41.7 41.7 50.8 31.5 30 58 63 45.2
-GS FFA 42.9 43.0 52.0 30.5 28.8 56.6 59.7 44.8
-GS cls token + WA 42.5 48.2 50.7 31.3 32.0 60.5 62.8 46.9
-GS FFA +WA 43.2 49.2 53.5 30.5 29.4 62.4 64.3 47.5
+Additional ablation studies are provided in the appendix.
+5 Discussions
+In this study, we introduce NIDS-Net, a framework designed for novel instance detection and segmentation.
+We utilize Grounding DINO and SAM to generate precise foreground object proposals, as opposed to conventional naive region proposals.
+We also present the Weight Adapter, which refines features from a pre-trained vision model effectively and mitigates the risk of overfitting.
+With the adapter, template and proposal embeddings of different instances are separated to facilitate the subsequent matching.
+Our method surpasses other approaches significantly in detection performance and also excels in segmentation compared to existing methods.
+However, there are limitations: given that our approach incorporates multiple pre-trained models, it requires greater computational resources compared to end-to-end detectors.
+Additionally, Grounding DINO may produce large bounding boxes that contain multiple objects, leading to detection failures.
+Moreover, our method tends to overlook heavily occluded objects with low confidence scores.
+In this work, instances are represented by $K$ template embeddings.
+For future research, we will explore using a single, distinctive embedding for each instance that acts as its identifier to enable one-shot detection.
+This approach will allow the detector to identify and locate a target instance within a query image using just one template image.
+Additionally, developing a computationally efficient method for this process will be crucial for its application in robotics.
+Acknowledgments
+This work was supported in part by the DARPA Perceptually-enabled Task Guidance (PTG) Program under contract number HR00112220005.
+Supplementary Material
+6 Training Details
+Detection.
+For detection datasets, the weight adapter is trained with a batch size of 1024, while the CLIP-Adapter is trained with a batch size of 512 to enhance performance.
+Both adapters are trained for the same number of epochs, as detailed in Table 6.
+To utilize Grounding DINO, a box threshold of 0.15 is set for the high-resolution and YCB-V datasets, and 0.10 for other datasets.
+Table 6: The training epochs of different detection datasets.
+High-resolution RoboTools LM-O YCB-V
+Both Adapters 40 80 40 40
+Segmentation.
+We combine all instances from seven core datasets of the BOP benchmark.
+We train both adapters with a batch size of 1344 (32 instances $\times$ 42 templates per instance) for 500 epochs.
+For Grounding DINO, a box threshold of 0.10 is set for all datasets.
+7 More Ablation Study
+FFA vs $cls$ token and Grounded SAM (GS) vs SAM in Detection Tasks.
+The detection capabilities of FFA and the $cls$ token are demonstrated using the high-resolution dataset, as presented in Table 9.
+For instance embedding generation, FFA achieves significantly greater enhancements compared to the $cls$ token.
+Furthermore, Table 9 illustrates that GS is effective in generating precise object proposals.
 Table 7: Detection results using different ViT backbones of Dinov2.
@@
 ViT-L/14 reg +WA 63.9 76.6 70.6
-Dinov2 backbones with adapter Our Weight Adapter is compatible with various backbones of Dinov2.
+Table 8: Detection results on the High-resolution real world detection dataset using various SAM variants.
+SAM Variant AP AP50 AP75 Time (sec)
+Mobile SAM (Tiny ViT) [97] 54.5 70.5 62.0 6.80
+Mobile SAM (Tiny ViT) [97] + WA 58.2 75.0 66.5 6.73
+HQ-SAM (ViT-H) [41] 59.8 71.5 65.7 7.38
+HQ-SAM (ViT H) [41] + WA 63.8 76.2 70.5 7.31
+SAM (ViT-H) (Ours) [47] 59.3 71.1 65.1 6.92
+SAM (ViT-H) + WA (Ours) [47] 63.9 76.6 70.6 6.78
+Table 9: The detection results on all images of the high-resolution dataset. “Proposal” refers to the object proposal method. “Embedding” denotes the method of instance embedding generation.
+Proposal Embedding AP AP50 AP75
+SAM $cls$ token 41.6 49.1 46.0
+GS $cls$ token 54.9 65.4 60.1
+GS FFA 59.3 71.1 65.1
+Image encoder.
+Given the same object proposals with GS, we evaluate the FFA embeddings from different image encoders on the High-resolution dataset [81].
+As illustrated in Table 10, DINOv2 exhibits superior performance attributable to its robust visual features.
+Fig.
+6 presents the visual results of various image encoders.
+Table 10: The instance embeddings of different image encoders.
+Image Encoder Feature Dimension AP AP50 AP75
+SAM (Vit-L) [47] 256 0.9 1.1 0.9
+SAM (Vit-L) [47] + WA 256 0.7 0.8 0.7
+DeiT III [86] (Vit-L) 784 2.9 3.6 3.1
+DeiT III [86] (Vit-L) + WA 784 2.8 3.4 3.0
+CLIP (Vit-L) [75] 1024 17.9 21.1 18.9
+CLIP (Vit-L) [75] + WA 1024 20.4 24.0 21.8
+DINOv2 (Vit-L) [68] 1024 56.8 67.7 62.3
+DINOv2 (Vit-L) [68] + WA 1024 60.4 71.8 66.4
+DINOv2 (Vit-L reg) [68] 1024 59.3 71.1 65.1
+DINOv2 (Vit-L reg) [68] + WA 1024 63.9 76.6 70.6
+Figure 6: Visual detection results on the High-resolution dataset using different image encoders.
+Dinov2 backbones with adapter.
+Our Weight Adapter is compatible with various backbones of Dinov2.
 Notably, more powerful backbones, which offer a more effective feature space, enable our adapter to deliver greater improvements.
@@
 Please refer to the Appendix for additional ablation studies.
-5 Discussions
-In this study, we introduce NIDS-Net, a framework designed for novel instance detection and segmentation.
-We utilize Grounding DINO and SAM to generate precise foreground object proposals, as opposed to conventional naive region proposals.
-We also present the Weight Adapter, which refines features from a pre-trained vision model effectively and mitigates the risk of overfitting.
-With the adapter, template and proposal embeddings of different instances are separated to facilitate the subsequent matching.
-Our method surpasses other approaches significantly in detection performance and also excels in segmentation compared to existing methods.
-However, there are limitations: given that our approach incorporates multiple pre-trained models, it requires a longer running time and greater computational resources compared to end-to-end detectors.
-Additionally, Grounding DINO may produce large bounding boxes that encompass multiple objects, leading to under-segmentation and detection failures.
-Moreover, our method tends to overlook heavily occluded objects with low confidence scores.
-In this work, instances are represented by $K$ template embeddings.
-For future research, we propose that each instance should be characterized by a single, distinctive embedding that acts as its identifier to enable one-shot detection.
-This approach will allow the detector to identify and locate a target instance within a query image using just one template image.
-Additionally, developing an efficient method for this process will be crucial for its application in robotics.
-Acknowledgments
-This work was supported in part by the DARPA Perceptually-enabled Task Guidance (PTG) Program under contract number HR00112220005.
+Aggregation.
+For cls token embeddings, averaging the top $k$ highest (avg k) scores yields the best results [66, 53].
+For FFA embeddings, the $\max$ aggregation function achieves optimal outcomes.
+The comparison of these aggregation functions is detailed in Table 11.
+SAM variants.
+We evaluate different SAM variants on the High-resolution real world detection dataset.
+As illustrated in Table 8, SAM and our weight adapter achieves the superior performance.
+Given the precise bounding boxes, HQ-SAM[41] yields results comparable to those of SAM [47].
+Table 11: Comparison of aggregation functions for segmentation performance.
+We report Average Precision (AP).
+avg k refers to averaging the top $k$ scores.
+All results are based on object proposals from Grounded SAM (GS).
+Embedding Aggregation BOP Datasets
+LM-O T-LESS TUD-L IC-BIN ITODD HB YCB-V Mean
+$cls$ token avg k 41.7 41.7 50.8 31.5 30 58 63 45.2
+$cls$ token $\max$ 42 37.4 45.9 30.2 28.9 54.9 61.5 43.0
+FFA avg k 42.5 42 47.4 28.2 27.3 55.1 61.5 43.4
+FFA $\max$ 42.9 43 52 30.5 28.8 56.6 59.7 44.8
+Runtime.
+We compare the efficiency of existing methods for novel instance segmentation, as presented in Table 12.
+Our approach significantly reduces running time by proposing only high-quality bounding boxes.
+Table 12: Runtime comparisons of various methods for novel instance segmentation.
+Method Proposal Server Time (sec)
+CNOS FastSAM Tesla V100 0.22
+CNOS DeForce RTX 3090 0.23
+SAM-6D (RGB) DeForce RTX 3090 0.25
+SAM-6D (RGBD) DeForce RTX 3090 0.45
+CNOS SAM Tesla V100 1.84
+CNOS DeForce RTX 3090 2.35
+SAM-6D (RGB) DeForce RTX 3090 2.28
+SAM-6D (RGBD) DeForce RTX 3090 2.80
+NIDS-Net (Ours) GS RTX A5000 0.49
+NIDS-Net + CA (Ours) 0.48
+NIDS-Net + WA (Ours) 0.48
+NIDS-Net + WA + $s_{\text{appe}}$ (Ours) 0.48
+8 Unseen Detection of BOP Benchmark
+We compare our approach with ZeroPose [13], CNOS [66], and SAM-6D [53] for 2D unseen detection, as illustrated in Table 13.
+Our method outperforms the best RGB method by 2.5 AP and competes effectively with the top RGB-D method.
+Table 13: Unseen instance detection results across the seven core datasets of the BOP benchmark, with all results reported as Average Precision (AP).
+Method Proposal BOP Datasets
+LM-O T-LESS TUD-L IC-BIN ITODD HB YCB-V Mean
+ZeroPose [13] SAM 36.7 30.0 43.1 22.8 25.0 39.8 41.6 34.1
+CNOS [66] SAM 39.5 33.0 36.8 20.7 31.3 42.3 49.0 36.1
+CNOS [66] FastSAM 43.3 39.5 53.4 22.6 32.5 51.7 56.8 42.8
+SAM-6D (RGB) [53] FastSAM 43.8 41.7 54.6 23.4 37.4 52.3 57.3 44.4
+SAM-6D (RGBD) [53] SAM 46.6 43.7 53.7 26.1 39.3 53.1 51.9 44.9
+SAM-6D (RGBD) [53] FastSAM 46.3 45.8 57.3 24.5 41.9 55.1 58.9 47.1
+NIDS-Net (Ours) GS 44.9 42.8 43.4 24.4 34.9 54.8 56.5 43.1
+NIDS-Net + WA(Ours) 44.9 48.9 46.0 24.5 36.0 59.4 62.4 46.0
+NIDS-Net + WA (Ours) + $s_{\text{appe}}$ 45.7 49.3 48.6 25.7 37.9 58.7 62.1 46.9
+9 More Qualitative Results
+9.1 Adapter
+To facilitate a comparison between the CLIP-Adapter and our Weight Adapter, we present a visual illustration in Figure 7.
+It is evident that the CLIP-Adapter alters the feature space and spoils the embeddings of background objects due to overfitting.
+In contrast, our Weight Adapter delivers robust embeddings within the original feature space.
+9.2 Detection
+We display the visual outcomes of our methodology employing the weight adapter on the LMO and YCB-V datasets in Figure 8.
+The gap between synthetic and real images results in some instances of detection failure.
+For example, some LM-O instances are not found with our method.
+9.3 Segmentation
+We present the visual outcomes of our approach using the weight adapter on the BoP datasets in Figures 9, 10, 11, and 12.
+These images demonstrate the effectiveness of our approach in cluttered environments.
+In some cases of T-LESS and IC-BIN datasets, Grounding DINO generates large bounding boxes which include multiple objects, causing under-segmentation.
+Furthermore, in IC-BIN and HB datasets, some heavily occluded objects with low confidence scores are overlooked by our method.
+Figure 7: Comparison of different adapters on a hard scene image from the high-resolution dataset. Red arrows denote background objects that are erroneously classified as targets.
+Figure 8: Visual detection results using the Weight Adapter on the LM-O and YCB-V datasets. Our approach detects object instances according to their synthetic template images.
+Figure 9: Qualitative segmentation results on the LM-O and T-Less datasets.
+Figure 10: Qualitative segmentation results on the TUD-L and IC-BIN datasets.
+Figure 11: Qualitative segmentation results on the ITODD and HB datasets
+Figure 12: Qualitative segmentation results on the YCB-V dataset.

### v3

@@
 Novel Instance Detection and Segmentation (NIDS) aims at detecting and segmenting novel object instances given a few examples of each instance.
-We propose a unified, simple yet effective framework (NIDS-Net) comprising object proposal generation, embedding creation for both instance templates and proposal regions, and embedding matching for instance label assignment.
+We propose a unified, simple, yet effective framework (NIDS-Net) comprising object proposal generation, embedding creation for both instance templates and proposal regions, and embedding matching for instance label assignment.
 Leveraging recent advancements in large vision methods, we utilize Grounding DINO and Segment Anything Model (SAM) to obtain object proposals with accurate bounding boxes and masks.
 Central to our approach is the generation of high-quality instance embeddings.
-We utilize foreground feature averages of patch embeddings from the DINOv2 ViT backbone, followed by refinement through a weight adapter mechanism that we introduce.
+We utilized foreground feature averages of patch embeddings from the DINOv2 ViT backbone, followed by refinement through a weight adapter mechanism that we introduce.
 We show experimentally that our weight adapter can adjust the embeddings locally within their feature space and effectively limit overfitting in the few-shot setting.
-This methodology enables a straightforward matching strategy, resulting in significant performance gains.
-Our framework surpasses current state-of-the-art methods, demonstrating notable improvements of 22.3, 46.2, 10.3, and 24.0 in average precision (AP) across four detection datasets.
-In instance segmentation tasks on seven core datasets of the BOP challenge, our method is around 4.5 times faster than the leading published RGB method and surpasses it by 3.6 AP.
-NIDS-Net is about 5.7 times faster than the top RGB-D method while maintaining competitive performance [^1] Project Page: https://irvlutd.github.io/NIDSNet/.
-1 Introduction
+Furthermore, the weight adapter optimizes weights to enhance the distinctiveness of instance embeddings during similarity computation.
+This methodology enables a straightforward matching strategy that results in significant performance gains.
+Our framework surpasses current state-of-the-art methods, demonstrating notable improvements in four detection datasets.
+In the segmentation tasks on seven core datasets of the BOP challenge, our method outperforms the leading published RGB methods and remains competitive with the best RGB-D method.
+We have also verified our method using real-world images from a Fetch robot and a RealSense camera.
+[^1] Project page with video, code and appendix: https://irvlutd.github.io/NIDSNet/
+I Introduction
+Novel Instance Detection and Segmentation (NIDS) is a crucial task in robot perception aimed at identifying and locating unseen instances in images or videos, given a few examples of each instance.
+Suppose that a robot needs to grasp a specific, novel object instance from a cluttered desk given only a small number of multi-view template images of the object.
+NIDS can provide the precise bounding box and mask of the target given a query image.
+The current paradigm for NIDS typically encompasses the following steps: (1) generating proposals from a query image, (2) obtaining embeddings of the proposals and the object templates, and (3) matching the embeddings of the proposals with those of the templates for identification.
+Recent work [1, 2, 3, 4] has utilized various open-world detectors, e.g., the Segment Anything Model (SAM) [5] or FastSAM [6], to obtain object proposals.
+However, the use of open-world detectors often results in the generation of region-based proposals rather than actual object proposals.
+This misidentification of regions as objects can lead to object identification issues.
+For example, a single object may be divided into multiple proposals, or some background regions may be misclassified as foreground objects.
+These false alarms cause issues for the following identification stage.
+To embed proposals and templates, some existing works, e.g., [2, 3, 4], adopt the $cls$ token of DINOv2 [7, 8].
+[1] used a 3D voxel representation.
+Ideally, for each specific unseen instance, the template embeddings from different views should be similar to each other, but markedly different from the embeddings of other instances.
+However, in these previous works, the embeddings of different instances may remain similar.
 Figure 1: We leverage pre-trained vision models for object proposal generation and feature extraction, and introduce a weight adapter to improve pre-trained feature embeddings for novel object instance detection and segmentation.
-Novel Instance Detection and Segmentation (NIDS) is a crucial task in computer vision aimed at identifying and locating unseen instances in images or videos, given a few examples of each instance.
-It also has wide applications in robotics.
-For example, suppose a robot needs to grasp a specific, novel object instance from a cluttered desk given only a small number of multi-view template images of the object.
-Ideally, detectors for this task should provide precise bounding boxes and masks of the target instances given a query image.
-However, standard detection methods trained on specific classes often struggle to recognize new objects.
-The current paradigm for NIDS typically encompasses the following steps: (1) generating proposals from a query image, (2) obtaining embeddings of the proposals and the object templates, and (3) matching the embeddings of proposals with those of the templates for identification.
-Recent work [49, 81, 66, 53] has utilized various open-world detectors to obtain object proposals.
-However, the use of models such as the Segment Anything Model (SAM) [47] or FastSAM [103] often results in the generation of region-based proposals rather than actual object proposals.
-This misidentification of regions as objects can lead to object identification issues.
-For example, a single object may be divided into multiple proposals or some background regions may be mistaken for foreground objects.
-These false alarms not only slow down the whole detection progress but also spoil the final results.
-To embed proposals and templates, some existing works, e.g., [81, 66, 53], adopt the $cls$ token of DINOv2  [68, 20] while others, e.g., Li et al.
-[49], employ a 3D voxel representation.
-Ideally, for each specific unseen instance, the template embeddings from different views should be similar to each other but markedly different from the embeddings of other instances.
-Despite these efforts, the embeddings of different instances may remain similar, and so the challenges persist.
-To address the limitations of the existing methods, we propose a framework for NIDS, as depicted in Fig.
+To address the limitations of existing methods, we propose a framework for NIDS, as depicted in Fig.
 1.
-Initially, we utilize Grounding DINO [56] with the text prompt “objects” on a cluttered query image to obtain bounding boxes for foreground objects.
-This method capitalizes on the inherent objectness of Grounding DINO using specialized prompts rather than standard labels.
-Subsequently, we employ SAM to create masks within these bounding boxes, generating precise object proposals comprising both bounding boxes and masks.
+Initially, we utilize Grounding DINO [9] with the text prompt “objects” on a cluttered query image to obtain bounding boxes of foreground objects.
+This method capitalizes on the inherent objectness of Grounding DINO using specialized prompts.
+Subsequently, we employ SAM [5] to create masks within these bounding boxes, generating precise object proposals comprising both bounding boxes and masks.
 For instance embeddings, we first extract average foreground features [10] from the patch embeddings of the DINOv2 ViT backbone.
 We then apply adapters to enhance instance embeddings by clustering similar instances and distancing different ones.
-The adapters are trained using the template images only, where InfoNCE loss [67, 15] is applied to the feature embeddings after the adapters.
+The adapters are trained using the template images only, where the InfoNCE loss [11, 12] is applied to the feature embeddings after the adapters.
 To refine these embeddings, the CLIP-Adapter [13] introduces a residual vector added to the original embedding, which risks overfitting with a few training examples.
-This addition can spoil and destabilize the embeddings of background objects, disrupting the entire feature space.
-To mitigate this issue, we, instead, introduce a weight adapter that modifies the original embeddings by applying learned weights.
-Finally, we employ a straightforward matching approach [62], such as stable matching or using $argmax$, to assign instance labels to these proposals.
-Our approach has been validated on four detection datasets and seven segmentation datasets for NIDS.
-The framework significantly surpasses existing state-of-the-art methods: it demonstrates substantial increases in average precision (AP), with gains of 22.3, 46.2, 10.3, and 24.0 across these detection datasets.
-Moreover, in instance segmentation tasks on the seven core datasets of the BOP challenge [84], our method outperforms the best published RGB and RGB-D methods.
+This addition can spoil and destabilize the embeddings of non-target objects, disrupting the entire feature space and causing the framework to misclassify non-targets as target instances.
+To mitigate this issue, we instead introduce a novel weight adapter (WA) that modifies the original embeddings by applying learned weights.
+Since the matching process relies on the standard cosine similarity, all embedding dimensions are treated equally, which may not sufficiently distinguish instances.
+The weights learned from our weight adapter emphasize the most relevant embedding channels.
+This produces more distinctive instance representations and improves performance by enhancing the discriminative capacity of embeddings.
+Finally, after applying the weight adapter to the embeddings, we employ a straightforward matching approach [14], e.g., using stable matching or $argmax$, to assign instance labels to these proposals from the query image.
+Our approach has been validated on four detection datasets, seven segmentation datasets, and in the real world for NIDS.
+The framework significantly surpasses existing state-of-the-art methods.
+It demonstrates substantial increases in average precision (AP), with gains of 22.3, 46.2, 10.3, and 24.0 across these detection datasets.
+Moreover, our method outperforms the best published RGB and RGB-D methods on instance segmentation tasks over the seven core datasets of the BOP challenge [15].
 Our contributions are summarized as follows.
-  • We propose NIDS-Net, a unified framework for novel instance detection and segmentation that surpasses existing published methods across four detection datasets and seven segmentation datasets of BOP challenge.
-  • We leverage the objectness of the open-world detector Grounding DINO and region recognition of Segmentation Anything Model (SAM) to obtain object proposals.
-  • We introduce the Weight Adapter, a tool designed to refine embeddings within their feature space while discouraging overfitting.
-This approach yields increasingly significant improvements as the adapting of the feature space becomes more robust and stable.
-2 Related Work
+  • We propose NIDS-Net, a unified framework for novel instance detection and segmentation that adapts pre-trained vision models for this task.
+  • We utilize the objectness of the Grounding DINO model and region recognition of the Segmentation Anything Model (SAM) to obtain object proposals.
+  • We introduce the Weight Adapter, a tool designed to refine embeddings within their feature space while preventing overfitting.
+This approach yields significant improvements as the adaptation of the feature space becomes more robust and stable.
+II Related Work
 Pretrained Models.
 Large-scale pretraining has demonstrated utility across various downstream tasks.
-Works such as [32, 91, 59, 24, 58, 10, 68] emphasize large-scale image representation learning, while [74, 50] focus on vision-language pretraining.
-Additionally, [3, 4, 96] address depth prediction, and [72, 73, 6, 87] tackle large-scale language modeling.
-Object detection is addressed by [79, 77, 43, 44, 31, 56], while image segmentation is the focus of [33, 89, 46].
-Works such as [94, 52, 14, 76] explore text-to-3D generation, and [82, 1] address text-to-video synthesis.
-Finally, 6DoF pose estimation and tracking are tackled by [90, 21, 51, 88].
-DINOv2 [68, 20] offers robust features to represent unseen instances.
+Works such as [16, 17, 7] emphasize large-scale image representation learning.
+DINOv2 [7] offers robust features to represent unseen instances.
 These foundational models offer diverse capabilities, with the challenge lying in effectively leveraging their wealth of knowledge for specific use cases, such as novel instance detection.
 In our work, we leverage pre-trained vision models for NIDS.
-Open-World Object Detection (OWOD).
-OWOD tackles real-world scenarios where object detection algorithms are tasked not only with identifying known and labeled objects but also with effectively handling novel and unknown objects during inference.
-This sets OWOD apart from traditional object detection, which confines itself to detecting only predetermined object classes.
-A key challenge in OWOD lies in detecting and classifying unknown objects that were not part of the training data.
-Recent approaches such as OWL-ViT [65], DITO [43], RO-ViT [44], YOLO-World [18], Sambor [31], and Grounding DINO [56] offer diverse solutions to address OWOD challenges.
-For example, OWL-ViT leverages a standard Vision Transformer [24] architecture with minimal modifications, incorporating contrastive image-text pre-training and end-to-end detection fine-tuning.
-DITO introduces an open-vocabulary detection approach based on detection-oriented image-text pretraining, replacing the classification architecture with a detector architecture during pretraining to better serve region-level recognition needs.
-Grounding DINO [56] utilizes features from DINO [10] to ground text queries.
-We employ Grounding DINO due to its strong performance.
-Similar to our work, Fang et al.
-[28] obtain object bounding boxes using the text prompt “a rigid object” with Grounding DINO.
-Few-Shot Object Detection (FSOD).
-FSOD involves detecting occurrences of a novel class, unseen during training, given only a few template examples.
-For example, De-ViT [102] introduces a region-propagation mechanism for localization, transforming propagated region masks into bounding boxes using a learnable spatial integral layer.
-FS-DETR [7], the first few-shot detection transformer, extends DETR [9] to process arbitrary numbers of novel objects concurrently while supporting various examples from each class.
-Other approaches such as [27] prioritize salient representations through angle distance and magnitude dispersion across few-shot samples,  [93] leverages variational autoencoders for data augmentation, and  [61] learns geometry-aware features with an offline simplex equiangular tight frame classifier.
-Additionally,  [60] generates specific and representative prototypes for each query image, considering both support and query image features.
-Our method is a FSOD method based on region proposal generation and matching.
-Compared to previous FSOD methods, it is simple yet effective.
 Instance Detection.
-Instance detection involves identifying an unseen instance in a test image using corresponding templates.
-Some methods, e.g.,  [64, 2, 57], rely on pure 2D representations and matching techniques.
-Specifically, DTOID [64] employs global object attention and a local pose-specific branch for template-guided heatmap prediction.
+Instance detection identifies an unseen instance in a test image using corresponding templates.
+Some methods such as [18, 19, 20], rely on pure 2D representations and matching techniques.
 However, these methods may struggle with variations in 2D appearance caused by occlusion or pose changes.
 In contrast, VoxDet [1] utilizes explicit 3D knowledge from multi-view templates, providing geometry-invariant representations.
-We generate 2D robust instance embeddings from templates with DINOv2 [68, 20].
+We generate 2D robust instance embeddings from templates with DINOv2 [7].
 Adapters for Pre-trained Models.
 The application of adapters atop large pre-trained models has emerged as a prominent strategy, yielding significant improvements across various tasks.
-Previous works have leveraged adapters to enhance few-shot image classification tasks [29, 99, 70, 101, 83].
-Other works have demonstrated effectiveness in language model fine-tuning [37, 39, 100, 30].
-Additional works exhibit the promise of adapters in vision tasks including detection and segmentation [17, 16, 71, 92, 23].
+Previous works have leveraged adapters to enhance few-shot image classification tasks [13, 21, 22].
 For novel instance detection and segmentation, we propose the Weight Adapter to enhance performance.
@@
 For instance, the SE block lacks a ReLU layer prior to the sigmoid function.
-3 Method
-Novel Instance Detection and Segmentation (NIDS), also referred to as open-world detection and segmentation, is a task to locate and label unseen instances within a query image.
-We will assume that each of the $N$ target instances is represented by $K$ template images $I_{T}\in\mathbb{R}^{K\times 3\times W\times H}$ and their corresponding segmentation masks.
+III Method
+The NIDS task is to locate and label novel object instances within a query image, given a set of template images of these objects.
+We assume that each of the $N$ target instances is represented by $K$ template images $I_{T}\in\mathbb{R}^{K\times 3\times W\times H}$ and their corresponding segmentation masks.
 During inference, the output of each query image $I_{Q}\in\mathbb{R}^{3\times W\times H}$ provides the bounding boxes for detection and instance masks for segmentation of these instances.
-3.1 Instance Embedding Generation Stage
+III-A Instance Embedding Generation Stage
 In our approach, an instance embedding summarizes the pixels of an instance.
@@
 We employ FFA to generate all initial instance embeddings using the object templates.
-3.2 Object Proposal Stage
+Figure 2: In our framework NIDS-Net, only adapters are learnable, while other models are frozen. Instance IDs are the instance labels.
+III-B Object Proposal Stage
 The purpose of this stage is to acquire object proposals from a query image.
@@
 2 illustrates the process of obtaining proposal embeddings.
-Figure 2: In our framework NIDS-Net, only adapters are learnable, while other models are frozen. Instance IDs are the instance labels.
-3.3 Embedding Refinement via an Adapter
+III-C Embedding Refinement via an Adapter
 Most instance embeddings are well separated.
 However, some embeddings from distinct instances may cluster together.
-For example, background object embeddings might be similar to targets.
+For example, non-target object embeddings might be similar to targets.
 Some target instances may resemble each other.
@@
 We train these adapters using the InfoNCE loss [11, 12], aiming to bring the embeddings of the same instance closer together while separating the embeddings of different instances.
-This training only uses the template images.
+This training only uses the few-shot template images.
 CLIP-Adapter (CA).
@@
 To mitigate this issue, CLIP-Adapter integrates residual connections that dynamically merge the newly adapted knowledge with the foundational knowledge from the original CLIP backbone.
-Given an instance embedding $f$ and an $MLP$ (Multi-Layer Perceptron), the CLIP-Adapter modifies the embedding according to the formula $f_{c}=\alpha\times MLP(f)+(1-\alpha)\times f$.
-In this equation, $f_{c}$ represents the adapted embedding, which is a linear combination of the transformed embedding $MLP(f)$ and the original embedding $f$.
+As illusrated in Fig.
+3 (Left), given an instance embedding f and an MLP (Multi-Layer Perceptron), the CLIP-Adapter modifies the embedding according to the formula $\mathbf{f_{c}}=\alpha\times\text{MLP}(\mathbf{f})+(1-\alpha)\times\mathbf{f}$.
+In this equation, $\mathbf{f_{c}}$ represents the adapted embedding, which is a linear combination of the transformed embedding $\text{MLP}(\mathbf{f})$ and the original embedding $\mathbf{f}$.
 The residual ratio $\alpha$ is set as 0.6.
+Figure 3: (Left) CLIP-Adapter [13] (Right) Our introduced weight adapter
 Our Weight Adapter (WA).
-The CLIP-Adapter refines instance embeddings by adding a new feature vector.
-However, this new vector might not align with the feature space of original instance embeddings, which can lead to overfitting.
-Since only several hundred template embeddings of target instances are trained and refined in some datasets, the embeddings of background objects can be spoiled as the process disrupts the distances between target and background instance embeddings.
-Assuming the feature space of original embeddings is robust and effective, it is crucial to fine-tune the instance embeddings within this space by constraining the scope of the adapted embeddings.
-This refinement can be accomplished by simply applying weights to the original instance embeddings.
-Hence, we introduce our Weight Adapter, a compact MLP network depicted in Fig. 3. The operation of this adapter is described through the following equations:
-$w$ $=\text{sigmoid}(MLP(\beta f)),$ (1)
-$f_{w}$ $=w\odot(\beta f).$ (2)
-Here, $\beta$ is a positive scalar.
-We set $\beta$ as 10.
-$w$ is the weights computed by passing the scaled embedding $\beta f$ through the MLP and then applying a $sigmoid$ function.
-The output $f_{w}$ is the element-wise product of $w$ and the scaled embedding $\beta f$, resulting in the weighted embedding.
-We include a ReLU layer before the $sigmoid$ function, ensuring that the sigmoid’s output is constrained to the range $[0.5,1)$.
-Subsequently, each refined embedding is close to its original counterpart.
-Additionally, $\beta$ allows for broader adjustments within the feature space.
-Given that cosine similarity is utilized to compare embeddings in the subsequent phase, our adapter alters the interrelationships among the values within each embedding.
-Further adapter details are presented in Fig.
-3.
-Figure 3: Comparison of two adapters. (left) CLIP-Adapter. (right) Our weight adapter.
-3.4 Matching Stage
+CLIP-Adapter refines instance embeddings by adding a new feature vector.
+However, this vector may not align with the original feature space, potentially causing overfitting.
+In datasets where only a few hundred template embeddings of target instances are refined, this adaptation can spoil non-target object embeddings by altering the relative distances between target and non-target instances.
+It leads the framework to misclassify non-targets as targets.
+Given the robustness and effectiveness of the original embedding space, it is essential to fine-tune instance embeddings within this space by constraining the adaptation.
+This can be achieved by applying weights to the original embeddings.
+We propose the Weight Adapter, a compact MLP-based network structure illustrated in Fig. 3 (Right). This adapter operates according to the following equations:
+$\mathbf{w}$ $=\text{sigmoid}(\text{MLP}(\beta\mathbf{f})),$ (1)
+$\mathbf{f_{w}}$ $=\mathbf{w}\odot(\beta\mathbf{f}),$ (2)
+where $\beta$ is a predefined positive scalar.
+$\mathbf{w}$ represents the weights derived by passing the scaled embedding $\beta\mathbf{f}$ through an MLP, followed by a sigmoid activation function.
+The resultant $\mathbf{f_{w}}$ is the element-wise multiplication of $\mathbf{w}$ and $\beta\mathbf{f}$, yielding the weighted embedding.
+An ReLU activation precedes the sigmoid in our architecture, restricting the output to the range [0.5, 1), which ensures that each adapted embedding retains proximity to its original value.
+We will use the cosine similarity between two embeddings to measure their similarity. For two embeddings $\mathbf{q}$ and $\mathbf{k}$, their cosine similarity is defined as:
+$\cos(\mathbf{q},\mathbf{k})=\frac{\mathbf{q}\cdot\mathbf{k}}{\|\mathbf{q}\|\;\|\mathbf{k}\|}.$ (3)
+With our weight adapter, we modify the embeddings using respective weight vectors $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$ as follows:
+$\mathbf{q}^{\prime}=\mathbf{w}_{1}\odot(\beta\mathbf{q}),$ (4)
+$\quad\mathbf{k}^{\prime}=\mathbf{w}_{2}\odot(\beta\mathbf{k}).$ (5)
+Consequently, the cosine similarity between $\mathbf{q}^{\prime}$ and $\mathbf{k}^{\prime}$ is expressed as:
+$\cos(\mathbf{q}^{\prime},\mathbf{k}^{\prime})=\frac{\sum_{i}w_{1,i}\,w_{2,i}\,q_{i}\,k_{i}}{\sqrt{\sum_{i}w_{1,i}^{2}q_{i}^{2}}\;\sqrt{\sum_{i}w_{2,i}^{2}k_{i}^{2}}}.$ (6)
+Here, $\beta$ acts as a scaling factor since the embedding values can be small.
+We simply set $\beta$ to 10.
+It facilitates feature scaling, stabilizing gradients and enhancing convergence.
+The computed weights from the adapter, $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, ensure that embeddings of the same instance are closely aligned.
+These weights allow us to find the important dimensions for similarity computation.
+Moreover, our adapter is versatile and flexible because it can be integrated with any image encoder or embedding generation mechanism.
+III-D Matching Stage
 This stage provides each proposal $i$ with an instance ID $o_{i}$ as its label and a confidence score $s_{i}$.
@@
 Each is defined by $\{b_{i},M_{i},o_{i},s_{i}\}$.
-Here, $b_{i}$ is the bounding box of an instance $M_{i}$ is the modal mask, which covers the visible object surface [84], $o_{i}$ is the instance ID, and $s_{i}$ is the confidence score.
+Here, $b_{i}$ is the bounding box of an instance.
+$M_{i}$ is the modal mask, which covers the visible object surface [15].
+$o_{i}$ is the instance ID, and $s_{i}$ is the confidence score.
 A confidence score threshold $\delta$ can be used to remove incorrect proposals.
-Table 1: Detection results on the high-resolution dataset.
+TABLE I: Detection results on the high-resolution dataset.
 avg indicates that we are evaluating all images, including those from the easy and hard scenes.
@@
 SAM + DINOv2f [7] 41.6 28.0 47.6 14.6 45.8 69.1 49.1 46.0 14.70
-NIDS-Net (Ours) 59.3 41.7 67.4 12.7 58.1 78.8 71.1 65.1 6.92
+NIDS-Net w/o adapter (Ours) 59.3 41.7 67.4 12.7 58.1 78.8 71.1 65.1 6.92
 NIDS-Net + CA (Ours) 61.3 42.5 69.6 17.0 59.3 81.6 73.2 67.0 6.99
 NIDS-Net + WA (Ours) 63.9 43.4 72.7 18.1 62.5 84.0 76.6 70.6 6.78
-4 Experiments
+IV Experiments
 We employ the Grounding DINO model with a Swin-T backbone [9] and a text prompt “objects”, and the default ViT-H SAM [5] to generate object proposals.
@@
 The experiments are conducted on an RTX A5000 GPU.
-For each dataset, we train both adapters on template embeddings using InfoNCE loss [67, 15] with temperature 0.05.
+For each dataset, we train both adapters on template embeddings using the InfoNCE loss [11, 12].
 With two linear layers, both adapters initially decrease the embedding dimension from $C$ to $C/4$, subsequently restoring it to $C$.
@@
 For additional details, please see the appendix.
-4.1 Detection Datasets
+IV-A Detection Datasets
 We utilize recently developed instance detection datasets and their associated baselines [1, 2].
@@
 High-resolution Dataset.
-Shen et al.
 [2] design a high-resolution dataset for instance detection.
@@
 Easy tags are assigned when objects are sparsely placed, while hard tags are used for more cluttered setups.
-According to Shen et al.
-[81], two types of baselines are set for the high-resolution dataset: Cut-Paste-Learn and a non-learned method.
+According to [2], two types of baselines are set for the high-resolution dataset: Cut-Paste-Learn and a non-learned method.
 Cut-Paste-Learn generates synthetic training images with 2D-box annotations, putting foreground instances in various sizes and aspect ratios onto different backgrounds.
@@
 Synthetic-Real Test Sets.
-Li et al.
 [1] employ two benchmarks for evaluation.
-LineMod-Occlusion (LM-O) [5] includes 8 texture-less objects and 1,514 bounding boxes, primarily challenging due to significant occlusion.
-The YCB-Video (YCB-V) [8] features 21 objects and 4,125 bounding boxes, with pose variation as the main challenge.
-Since these datasets have real testing images without reference videos, Li et al.
-[49] generate synthetic videos using CAD models.
+LineMod-Occlusion (LM-O) [33] includes 8 texture-less objects and 1,514 bounding boxes.
+The YCB-Video (YCB-V) [34] features 21 objects and 4,125 bounding boxes.
+Since these datasets have real testing images without reference videos, [1] generate synthetic videos using CAD models.
 We sample 16 synthetic template images per object from these videos.
-RoboTools Benchmark.
-Li et al.
-[49] introduce RoboTools, a challenging real-world benchmark for robotics.
-It features 20 distinct instances, 9,109 annotations, and 24 complex scenarios.
-Compared to previous benchmarks [5, 8], RoboTools offers greater difficulty due to its cluttered backgrounds and significant pose variations.
-Additionally, the benchmark’s reference videos are made from real images and incorporate realistic lighting conditions, including shadows.
-25 real template images of each instance are sampled from reference videos.
-For cluttered scenes from the datasets RoboTools, LM-O and YCB-V, Li et al.
-[49] have developed several 2D baselines: OLNDINO, OLNCLIP, and OLNCorr.
+RoboTools Benchmark [1] features 20 distinct instances, 9,109 annotations, and 24 complex scenarios.
+25 real template images per instance are sampled from their reference videos.
+For cluttered scenes from the datasets RoboTools, LM-O and YCB-V, [1] have developed several 2D baselines: OLNDINO, OLNCLIP, and OLNCorr.
 Initially, they generate open-world 2D proposals using their detection module [35].
@@
 They also employ the class-level one-shot detectors OS2D [37] and BHRL [38].
-To address the limitations of traditional 2D methods with pose variations and occlusions, Li et al.
-[49] introduced VoxDet which is based on a 3D voxel representation.
-To train these methods, Li et al.
-[49] developed the open-world instance detection (OWID-10k) dataset as a synthetic instance detection training set.
-It features 9,901 objects from ShapeNet [12] and ABO [19], each rendered into a 40-frame, 360-degree video using Blenderproc [22].
+To address the limitations of traditional 2D methods with pose variations and occlusions, [1] introduced VoxDet which is based on a 3D voxel representation.
+To train these methods, [1] developed a synthetic instance detection dataset (OWID-10k).
 Evaluation Metrics.
 For detection, we evaluate our method with Average Precision (AP).
-AP is computed by averaging precision scores at various Intersection over Union (IoU) thresholds, specifically from 0.5 to 0.95, in increments of 0.05.
-This metric is the primary evaluation standard in the COCO Object Detection dataset [54].
+AP is computed by averaging precision scores at various Intersection over Union (IoU) thresholds, specifically from 0.5 to 0.95, in increments of 0.05 [39].
 Additionally, AP50 and AP75 are variations of this metric, where precision is averaged across all instances at IoU thresholds of 0.5 and 0.75, respectively.
-4.2 Segmentation Datasets
+IV-B Segmentation Datasets
 The BOP Challenge.
 We test our method using seven datasets from the BOP challenge [15]: LineMod Occlusion (LM-O) [33], T-LESS [40], TUD-L [41], IC-BIN [42], ITODD [43], HomebrewedDB (HB) [44], and YCB-Video [34].
-The objects from these datasets vary widely.
 We use 42 template rendering images from CNOS [3], which are generated via BlenderProc [45].
@@
 The AP is computed by averaging the precision scores at various IoU thresholds, from 0.50 to 0.95, increasing by 0.05 at each step.
-Figure 4: Visual examples of our results using the Weight Adapter on the high-resolution and RoboTools datasets.
-Our approach detects specific object instances in cluttered scenes according to their real template images.
-Table 2: Detection performance on the fully real dataset, RoboTools.
+TABLE II: Detection performance on the fully real dataset, RoboTools.
 Proposal indicates the object proposal model.
@@
 VoxDet [1] OLN* 18.7 23.6 20.5
-NIDS-Net (Ours) GS 63.3 77.2 68.7
+NIDS-Net w/o adapter (Ours) GS 63.3 77.2 68.7
 NIDS-Net + CA (Ours) GS 64.8 79.2 70.7
 NIDS-Net + WA (Ours) GS 64.9 79.4 70.8
-Table 3: Detection performance on the LM-O and YCB-V datasets.
+TABLE III: Detection performance on the LM-O and YCB-V datasets.
 OLN* is trained alongside the matching head whereas OLN uses fixed modules.
@@
 VoxDet [1] OLN* OWID 29.2 43.1 33.3 31.5 51.3 33.4 30.4 47.2 33.4 0.154
-NIDS-Net (Ours) GS N/A 38.7 66.0 41.0 53.0 72.9 61.7 45.9 69.5 51.4 3.73
+NIDS-Net w/o adapter (Ours) GS N/A 38.7 66.0 41.0 53.0 72.9 61.7 45.9 69.5 51.4 3.73
 NIDS-Net + CA (Ours) GS N/A 39.2 67.0 41.4 53.9 74.1 62.7 46.6 70.6 52.1 3.71
 NIDS-Net + WA (Ours) GS N/A 39.5 67.4 41.8 55.5 75.5 65.0 47.5 71.5 53.4 3.73
-Table 4: Novel instance segmentation results on the seven core datasets of the BOP benchmark.
+TABLE IV: Novel instance segmentation results on the seven core datasets of the BOP benchmark.
 We utilize Average Precision (AP) to compare these methods.
@@
 SAM-6D (SAM, RGBD[4] 46.0 45.1 56.9 35.7 33.2 59.3 60.5 48.1
-NIDS-Net (Ours) 42.9 43.0 52.0 30.5 28.8 56.6 59.7 44.8
+NIDS-Net w/o adapter (Ours) 42.9 43.0 52.0 30.5 28.8 56.6 59.7 44.8
 NIDS-Net + CA (Ours) 42.8 47.5 50.0 29.7 27.6 62.7 63.5 46.3
@@
 NIDS-Net + WA + $s_{\text{appe}}$ (Ours) 43.9 49.6 55.6 32.8 31.5 62.0 65.0 48.6
-Table 5: Comparison of segmentation performance across various proposal models and embedding generation options. The first row is from CNOS [66].
+TABLE V: Segmentation performance comparison across proposal models and embedding methods. The first row is from CNOS [3].
 Proposal Embedding BOP Datasets
@@
 GS FFA + WA 43.2 49.2 53.5 30.5 29.4 62.4 64.3 47.5
-Figure 5: Comparison of segmentation results using CNOS, SAM6D, and our NIDS-Net on the YCB-V dataset.
-CNOS and SAM6D may misclassify some background regions or object parts as objects due to their reliance on SAM for proposal generation.
+Figure 4: Visual results on the RoboTools benchmark.
+Figure 5: Comparison of segmentation results using CNOS, SAM6D, and NIDS-Net on the YCB-V dataset.
+CNOS and SAM6D may misclassify some background regions or object parts as objects due to proposal generation of SAM.
+Red arrows indicate these mistakes.
 NIDS-Net addresses this limitation with Grounded-SAM.
-4.3 Benchmarking Results
+IV-C Benchmarking Results
 Detection.
-The high-resolution dataset results are presented in Table 1.
+The high-resolution dataset results are presented in Table I.
 Our models dramatically outperform existing techniques, achieving the highest AP scores across all categories.
@@
 Additionally, as this dataset contains 100 instances with some similarities among them, our Weight Adapter boosts overall performance by 4.6 AP and improves the detection of small objects by 5.4 AP.
-For the RoboTools dataset, detailed in Table 2, our method outperforms the state-of-the-art VoxDet by 46.2 AP.
+For the RoboTools dataset, detailed in Table II, our method outperforms the state-of-the-art VoxDet by 46.2 AP.
 Since RoboTools contains only 20 instances, our adapter has limited scope to enhance performance.
 Our approach achieves over 60 AP on these two fully real datasets.
-For the Synthetic-Real datasets, LM-O and YCB-V, we present a comparison of our approach with other methods in Table 3.
+For the Synthetic-Real datasets, LM-O and YCB-V, we present their results in Table III.
 Our model outperforms VoxDet by 10.3 AP on LM-O and by 24.0 AP on YCB-V.
@@
 In contrast, we generate proposals using Grounded-SAM and acquire embeddings through the FFA pipeline.
-The segmentation results for the seven principal datasets of the BOP challenge are detailed in Table 4.
+The segmentation results for the seven principal datasets of the BOP challenge are detailed in Table IV.
 Overall, our method surpasses the state-of-the-art SAM-6D and achieves superior results compared to RGBD results.
-Qualitative Result.
-We present detection results for both simple and complex scenes in Fig.
-4.
-Fig.
-5 displays a visual comparison of the segmentation results.
-Additional examples are available in the appendix.
+Qualitative Results. We present detection results of RoboTools in Fig. 4. Moreover, Fig. 5 displays a visual comparison of the segmentation results. Additional examples are available in the appendix.
 Runtime.
-The tables of detection results include the runtime of our method.
+Table I includes the runtime of our method for object detection.
 Grounded-SAM substantially decreases the number of proposals compared to SAM, thereby accelerating the entire process.
 Please see the appendix for segmentation runtime.
-4.4 Ablation Study
-Grounded SAM (GS) vs SAM.
-On the seven BOP datasets containing images of numerous cluttered scenes, we compare the object proposals from GS and SAM as detailed in Table 5.
+IV-D Real-world Testing and Failure Cases
+We capture 14 multi-view template images per object using seven Intel RealSense D455 cameras for real-world objects.
+With a trained weight adapter on these images, our method is tested on these objects with one D455 camera, demonstrating robust performance across various scenes (Fig.
+6).
+However, it misclassifies objects when highly similar instances are present as in Fig.
+7.
+With the weight adapter trained on BOP datasets, we also tested our method on YCB-V objects using a Fetch robot.
+The demonstrations are included in the supplementary video.
+Figure 6: Evaluation of NIDS-Net on real-world objects across diverse scenes. All predictions are accurate in these examples.
+Figure 7: Examples of failure cases from NIDS-Net, where the orange arrows indicate incorrect predictions.
+TABLE VI: The performance comparison of $\beta$ in the Weight Adapter on the high-resolution dataset.
+$\beta$ 1 5 10 20 50 100
+AP 62.0 63.1 63.9 63.1 62.7 62.2
+AP50 73.7 75.6 76.6 75.6 75.1 74.3
+AP75 68.3 69.5 70.6 69.3 69.0 68.2
+IV-E Ablation Study
+Grounded-SAM (GS) vs SAM.
+On the seven BOP datasets containing images of numerous cluttered scenes, we compare the object proposals from GS and SAM as detailed in Table V.
 The results indicate that GS yields more precise bounding boxes and masks compared to using SAM alone.
@@
 FFA vs $cls$ token.
-To compare these two types of embedding generation, we present the results in Table 5.
+To compare these two types of embedding generation, we present the results in Table V.
 Our weight adapter enhances both two types of embeddings.
 Despite close segmentation results, FFA produces embeddings that possess greater adaptive potential, demonstrated by higher AP scores after adaption via our weight adapter.
+$\beta$ in our Weight Adapter.
+We evaluate various values of $\beta$ on the high-resolution dataset.
+All results presented in Table VI surpass those obtained using CLIP-Adapter.
+We set $\beta$ to to 10 across all datasets as it yields the best results.
 Additional ablation studies are provided in the appendix.
-5 Discussions
+V Discussions
 In this study, we introduce NIDS-Net, a framework designed for novel instance detection and segmentation.
 We utilize Grounding DINO and SAM to generate precise foreground object proposals, as opposed to conventional naive region proposals.
-We also present the Weight Adapter, which refines features from a pre-trained vision model effectively and mitigates the risk of overfitting.
+We also introduce the Weight Adapter, which effectively refines features from a pre-trained vision model, mitigates the risk of overfitting, and adjusts the weighting of cosine similarities.
 With the adapter, template and proposal embeddings of different instances are separated to facilitate the subsequent matching.
 Our method surpasses other approaches significantly in detection performance and also excels in segmentation compared to existing methods.
-However, there are limitations: given that our approach incorporates multiple pre-trained models, it requires greater computational resources compared to end-to-end detectors.
-Additionally, Grounding DINO may produce large bounding boxes that contain multiple objects, leading to detection failures.
-Moreover, our method tends to overlook heavily occluded objects with low confidence scores.
+However, there are limitations.
+Given that our approach incorporates multiple pre-trained models, it requires greater computational resources compared to end-to-end detectors.
+When instances exhibit highly similar appearances, NIDS-Net may encounter detection failures.
+The method sometimes missed heavily occluded objects with low confidence scores.
 In this work, instances are represented by $K$ template embeddings.
@@
 Additionally, developing a computationally efficient method for this process will be crucial for its application in robotics.
-Acknowledgments
+ACKNOWLEDGMENT
 This work was supported in part by the DARPA Perceptually-enabled Task Guidance (PTG) Program under contract number HR00112220005.
-Supplementary Material
-6 Training Details
+APPENDIX
+VI Training Details
 Detection.
 For detection datasets, the weight adapter is trained with a batch size of 1024, while the CLIP-Adapter is trained with a batch size of 512 to enhance performance.
-Both adapters are trained for the same number of epochs, as detailed in Table 6.
+Both adapters are trained for the same number of epochs, as detailed in Table VII.
 To utilize Grounding DINO, a box threshold of 0.15 is set for the high-resolution and YCB-V datasets, and 0.10 for other datasets.
-Table 6: The training epochs of different detection datasets.
+TABLE VII: The training epochs of different detection datasets.
 High-resolution RoboTools LM-O YCB-V
@@
 For Grounding DINO, a box threshold of 0.10 is set for all datasets.
-7 More Ablation Study
-FFA vs $cls$ token and Grounded SAM (GS) vs SAM in Detection Tasks.
-The detection capabilities of FFA and the $cls$ token are demonstrated using the high-resolution dataset, as presented in Table 9.
-For instance embedding generation, FFA achieves significantly greater enhancements compared to the $cls$ token.
-Furthermore, Table 9 illustrates that GS is effective in generating precise object proposals.
-Table 7: Detection results using different ViT backbones of Dinov2.
+VII More Ablation Study
+TABLE VIII: Detection results using different ViT backbones of Dinov2.
 “reg” indicates DINOv2 with registers.
@@
 ViT-L/14 reg +WA 63.9 76.6 70.6
-Table 8: Detection results on the High-resolution real world detection dataset using various SAM variants.
+TABLE IX: Detection results on the High-resolution real-world detection dataset using various SAM variants.
 SAM Variant AP AP50 AP75 Time (sec)
@@
 SAM (ViT-H) + WA (Ours) [5] 63.9 76.6 70.6 6.78
-Table 9: The detection results on all images of the high-resolution dataset. “Proposal” refers to the object proposal method. “Embedding” denotes the method of instance embedding generation.
+TABLE X: The detection results on all images of the high-resolution dataset. “Proposal” refers to the object proposal method. “Embedding” denotes the method of instance embedding generation.
 Proposal Embedding AP AP50 AP75
@@
 Given the same object proposals with GS, we evaluate the FFA embeddings from different image encoders on the High-resolution dataset [2].
-As illustrated in Table 10, DINOv2 exhibits superior performance attributable to its robust visual features.
+As illustrated in Table XI, DINOv2 exhibits superior performance attributable to its robust visual features.
 Fig.
 8 presents the visual results of various image encoders.
-Table 10: The instance embeddings of different image encoders.
+TABLE XI: The instance embeddings of different image encoders.
 Image Encoder Feature Dimension AP AP50 AP75
@@
 Notably, more powerful backbones, which offer a more effective feature space, enable our adapter to deliver greater improvements.
-Details of this comparison are provided in Table 7.
-Please refer to the Appendix for additional ablation studies.
+Details of this comparison are provided in Table VIII.
 Aggregation.
@@
 For FFA embeddings, the $\max$ aggregation function achieves optimal outcomes.
-The comparison of these aggregation functions is detailed in Table 11.
+The comparison of these aggregation functions is detailed in Table XII.
 SAM variants.
 We evaluate different SAM variants on the High-resolution real world detection dataset.
-As illustrated in Table 8, SAM and our weight adapter achieves the superior performance.
+As illustrated in Table IX, SAM and our weight adapter achieves the superior performance.
 Given the precise bounding boxes, HQ-SAM[49] yields results comparable to those of SAM [5].
-Table 11: Comparison of aggregation functions for segmentation performance.
+TABLE XII: Comparison of aggregation functions for segmentation performance.
 We report Average Precision (AP).
@@
 Runtime.
-We compare the efficiency of existing methods for novel instance segmentation, as presented in Table 12.
+We compare the efficiency of existing methods for novel instance segmentation, as presented in Table XIII.
 Our approach significantly reduces running time by proposing only high-quality bounding boxes.
-Table 12: Runtime comparisons of various methods for novel instance segmentation.
+TABLE XIII: Runtime comparisons of various methods for novel instance segmentation.
 Method Proposal Server Time (sec)
-CNOS FastSAM Tesla V100 0.22
-CNOS DeForce RTX 3090 0.23
-SAM-6D (RGB) DeForce RTX 3090 0.25
-SAM-6D (RGBD) DeForce RTX 3090 0.45
-CNOS SAM Tesla V100 1.84
-CNOS DeForce RTX 3090 2.35
-SAM-6D (RGB) DeForce RTX 3090 2.28
-SAM-6D (RGBD) DeForce RTX 3090 2.80
-NIDS-Net (Ours) GS RTX A5000 0.49
+CNOS [3] FastSAM Tesla V100 0.22
+CNOS [3] DeForce RTX 3090 0.23
+SAM-6D (RGB) [4] DeForce RTX 3090 0.25
+SAM-6D (RGBD) [4] DeForce RTX 3090 0.45
+CNOS [3] SAM Tesla V100 1.84
+CNOS [3] DeForce RTX 3090 2.35
+SAM-6D (RGB) [4] DeForce RTX 3090 2.28
+SAM-6D (RGBD) [4] DeForce RTX 3090 2.80
+NIDS-Net w/o adapter (Ours) GS RTX A5000 0.49
 NIDS-Net + CA (Ours) 0.48
@@
 NIDS-Net + WA + $s_{\text{appe}}$ (Ours) 0.48
-8 Unseen Detection of BOP Benchmark
-We compare our approach with ZeroPose [13], CNOS [66], and SAM-6D [53] for 2D unseen detection, as illustrated in Table 13.
+VIII Unseen Detection of BOP Benchmark
+We compare our approach with ZeroPose [46], CNOS [3], and SAM-6D [4] for 2D unseen detection, as illustrated in Table XIV.
 Our method outperforms the best RGB method by 2.5 AP and competes effectively with the top RGB-D method.
-Table 13: Unseen instance detection results across the seven core datasets of the BOP benchmark, with all results reported as Average Precision (AP).
+TABLE XIV: Unseen instance detection results across the seven core datasets of the BOP benchmark, with all results reported as Average Precision (AP).
 Method Proposal BOP Datasets
@@
 SAM-6D (RGBD) [4] FastSAM 46.3 45.8 57.3 24.5 41.9 55.1 58.9 47.1
-NIDS-Net (Ours) GS 44.9 42.8 43.4 24.4 34.9 54.8 56.5 43.1
+NIDS-Net w/o adapter (Ours) GS 44.9 42.8 43.4 24.4 34.9 54.8 56.5 43.1
 NIDS-Net + WA (Ours) 44.9 48.9 46.0 24.5 36.0 59.4 62.4 46.0
-NIDS-Net + WA (Ours) + $s_{\text{appe}}$ 45.7 49.3 48.6 25.7 37.9 58.7 62.1 46.9
-9 More Qualitative Results
-9.1 Adapter
+NIDS-Net + WA + $s_{\text{appe}}$ (Ours) 45.7 49.3 48.6 25.7 37.9 58.7 62.1 46.9
+IX More Qualitative Results
+IX-A Adapter
 To facilitate a comparison between the CLIP-Adapter and our Weight Adapter, we present a visual illustration in Figure 9.
-It is evident that the CLIP-Adapter alters the feature space and spoils the embeddings of background objects due to overfitting.
+It is evident that the CLIP-Adapter alters the feature space and spoils the embeddings of non-target objects due to overfitting.
 In contrast, our Weight Adapter delivers robust embeddings within the original feature space.
-9.2 Detection
-We display the visual outcomes of our methodology employing the weight adapter on the LMO and YCB-V datasets in Figure 8.
+IX-B Detection
+We display the visual outcomes of our methodology with the weight adapter on the LMO and YCB-V datasets in Figure 11.
 The gap between synthetic and real images results in some instances of detection failure.
 For example, some LM-O instances are not found with our method.
-9.3 Segmentation
-We present the visual outcomes of our approach using the weight adapter on the BoP datasets in Figures 9, 10, 11, and 12.
+The examples of the high-resolution dataset are presented in Fig.
+10.
+IX-C Segmentation
+We present the visual results of our approach using the weight adapter on the BoP datasets in Figures 12, 13, 14, and 15.
 These images demonstrate the effectiveness of our approach in cluttered environments.
@@
 Furthermore, in IC-BIN and HB datasets, some heavily occluded objects with low confidence scores are overlooked by our method.
-Figure 7: Comparison of different adapters on a hard scene image from the high-resolution dataset. Red arrows denote background objects that are erroneously classified as targets.
+Figure 9: Comparison of different adapters on a hard scene image from the high-resolution dataset. Red arrows denote non-target objects that are erroneously classified as targets.
+Figure 10: Visual examples of our results using the Weight Adapter on the high-resolution dataset.
+Our approach detects specific object instances in cluttered scenes according to their real template images.
 Figure 11: Visual detection results using the Weight Adapter on the LM-O and YCB-V datasets. Our approach detects object instances according to their synthetic template images.